### v1

# Position: The Hidden Costs and Measurement Gaps of Reinforcement Learning with Verifiable Rewards

Abstract
Reinforcement learning with verifiable rewards (RLVR) is a practical and scalable approach to enhancing large language models in areas such as math, code, and other structured tasks. Two questions motivate this paper: how much of the reported gains survive under strictly parity-controlled evaluation, and whether RLVR is cost-free or exacts a measurable tax. We argue that progress is real, but gains are often overstated due to three forces—an RLVR tax, evaluation pitfalls, and data contamination. Using a partial-prompt contamination audit and matched-budget reproductions across base and RL models, we show that several headline gaps shrink or vanish under clean, parity-controlled evaluation. We then propose a tax-aware training and evaluation protocol that co-optimizes accuracy, grounding, and calibrated abstention and standardizes budgeting and provenance checks. Applied to recent RLVR setups, this protocol yields more reliable estimates of reasoning gains and, in several cases, revises prior conclusions. Our position is constructive: RLVR is valuable and industry-ready; we advocate keeping its practical benefits while prioritizing reliability, safety, and measurement.
1 Introduction
Reinforcement learning with verifiable rewards (RLVR) has become a leading post-training route for improving large language models on math, code, and other structured tasks  (Luong et al., 2024; Wen et al., 2025a). By optimizing against automatically computable signals—unit tests for programs, exact numeric or string matches for math, or retrieval-grounded checks for citations—RLVR promises a scalable, label-efficient path to better reasoning. Recent results are striking: across multiple domains, RLVR systems often post large gains on standard benchmarks. Moreover, Figure 2 shows a rise in RLVR-tagged papers alongside improvements on AIME-24/25 through 2024–H1 2025, underscoring both the field’s momentum and the need to separate genuine reasoning gains from measurement and budgeting artifacts.
Beneath the surface, however, sits a question that frames the current debate: does RLVR genuinely impart new reasoning capability, or does it mainly sharpen selection among behaviors the base model already knows how to produce? Parity-controlled studies show that base models can narrow or erase RLVR gaps when given matched sampling budgets—consistent with smarter search rather than capability expansion (Yue et al., 2025; Wu et al., 2025a). At the same time, several regimes report gains that are hard to recover by sampling alone: explicitly optimizing the multi-sample objective (pass@k training) (Wang et al., 2025b), curriculum via self-play with variational problem synthesis (Zhang et al., 2025a), and distribution-aware reward shaping that counters rank bias and diversity collapse (Liu et al., 2025b), alongside longer-horizon RL schedules and unlikeliness rewards (Liu et al., 2025a; He et al., 2025a). Even metric choice can flip conclusions: answer-only scores may diverge from process-aware metrics that require both a correct answer and a valid chain of thought.
A second complication is what we call the RLVR tax: unintended costs that can ride along with apparent gains. In practice, RLVR tends to reduce abstention and increase stated confidence—sometimes even when answers are wrong—thereby shifting risk from “I don’t know” to assertive errors (Song et al., 2025; Mei et al., 2025; Yao et al., 2025). It can also chip away at instruction fidelity on longer generations, where adhering to formats or constraints becomes harder as chains grow (Fu et al., 2025; Li et al., 2025). Finally, longer, more explicit reasoning traces expand the attack and leakage surface, raising jailbreak success and privacy exposure if left unchecked (Zhou et al., 2025; Jiang et al., 2025; Ackerman & Panickssery, 2025; Green et al., 2025; Huang et al., 2025; Zhang et al., 2025b).
A third, orthogonal complication is measurement. Reported advances are sensitive to sampling budgets and decoding settings, to the stability of LLM-as-a-judge evaluators, to calibration drift, and to data provenance (Hochlehnert et al., 2025; Chandak et al., 2025; Brown et al., 2024; Muennighoff et al., 2025; Zhao et al., 2025b; Wang et al., 2024a; b; Leng et al., 2024; Shen et al., 2025; Wu et al., 2025b; He et al., 2025b; Mirzadeh et al., 2024; Liu et al., 2024b). When budgets are matched, judges are stress-tested, and datasets are versioned and decontaminated, reported improvement gaps can diminish, suggesting that part of the apparent progress reflects evaluation design rather than durable capability (Hochlehnert et al., 2025; Chandak et al., 2025). To help orient the reader, Figure 1 summarizes three threads (taxes, evaluation pitfalls, and contamination) and previews the tax-aware protocol we adopt.
Figure 1: Paper Roadmap: taxes, evaluation pitfalls, contamination, and the unified protocol.
Scope and definitions.
Throughout, we use RLVR to mean post-training that optimizes LLMs against automatically checkable signals rather than human preference models (Ouyang et al., 2022). Typical implementations adopt PPO/GRPO/DAPO-style updates with a KL penalty to the base policy, often mixing offline rollouts with online sampling, adding entropy regularization, and using selective filtering or gating. Rewards are frequently componentized—for correctness, grounding/citation sufficiency, and calibrated refusal—and introduced in stages. Our analysis centers on open-weight models fine-tuned for math/code/question answering (QA) with verifiable objectives, in both single- and multi-domain regimes. We deliberately avoid a catalogue of low-level algorithmic variants except where they bear on reliability (e.g., unlikeliness rewards that counter rank bias, or curricula that change variance). The primary threats considered are privacy leakage from long chains of thought (CoT) and increased jailbreak susceptibility during or after RLVR.
Figure 2: Monthly RLVR activity vs. AIME performance (time span: May 2024–June 2025). Left axis: count of pages per month whose title or abstract contains “RLVR” or “reinforcement learning with verifiable rewards” (Google Scholar and arXiv via SerpAPI). Right axis: pass@1 (%) on AIME-24 (circles, warm) and AIME-25 (squares, cool) for selected models; labels show model names.
Position and Roadmap.
RLVR is valuable and industry-ready, but we encourage prioritizing reliability, safety, and measurement. We (i) synthesize evidence on sharpening vs. expansion (section 2); (ii) analyze the RLVR tax and how it distorts reported improvements (section 3); (iii) reproduce representative gaps under parity controls (section 4); (iv) audit contamination with partial-prompt probes (section 5); and (v) offer a tax-aware protocol that co-optimizes accuracy, grounding, and calibrated abstention while standardizing budgeting and provenance checks (section 6).
2 What Changes Under RLVR? An Evidence Overview
We review empirical signals about RLVR’s effects without taking a side on the “expansion v.s. sharpening” question. Skeptical and optimistic perspectives recur in the literature, and each tends to appear under identifiable conditions.
2.1 Skeptical Lens: Sharpening and Sampling Effects
Under a skeptical lens, RLVR often appears to improve sample efficiency and steer the model toward high-reward regions already present in the base distribution, rather than broaden fundamental capability (Gandhi et al., 2025; Shah et al., 2025). When base models are evaluated under matched sampling budgets (i.e., pass@k), gaps between base and RLVR-trained models often shrink (Yue et al., 2025), consistent with selection rather than acquisition. Empirically, outputs can collapse toward a dominant pretraining mode with reduced diversity (Zhao et al., 2025a), and theory suggests that after RLVR, support shrinkage tends to outweigh support expansion (Wu et al., 2025a). Stress tests add caution: performance collapses at higher complexities have been reported even with ample compute (Shojaee et al., 2025), although some failures trace to evaluation artifacts such as token limits and unsolvable items, underscoring the need for budget parity, seed control, and dataset hygiene (see Section 4(Opus & Lawsen, 2025). Finally, several works separate knowledge injection from policy optimization: pipelines that use SFT or distillation to inject knowledge and RLVR to select among candidates typically beat pure GRPO, suggesting RLVR optimizes within learned support (Ma et al., 2025; Wan et al., 2025; Chen et al., 2025a); guidance and self-distillation further help over vanilla GRPO (Nath et al., 2025; Liu et al., 2025c).
2.2 Optimistic Lens: Signals of Capability Broadening
Under an optimistic lens, carefully designed RLVR appears to expand what models can do, not merely sharpen selection. Prolonged optimization with explicit KL control and periodic resets (ProRL) reports trajectories that are not reachable by sampling the base policy alone, with the largest gains where the base initially struggles (Liu et al., 2025a). Reward shaping also matters: by diagnosing a GRPO tendency to reinforce high-probability completions, He et al. (2025a) show that adding unlikeliness rewards and training deeper widens the model’s support and improves multi-sample theorem proving, indicating qualitative changes to reachable reasoning modes. These effects are conditioned by pretraining coverage: cross-domain analyses find broader benefits in Math/Code/Science, while under-represented areas such as Logic/Simulation/Tabular realize meaningful gains when RLVR is applied in-domain (Cheng et al., 2025b). Consistent with this, process-aware metrics that require both a correct answer and a syntactically valid chain (CoT-pass@k) often reveal larger RLVR deltas than answer-only measures (Wen et al., 2025b). While generalization to substantially harder regimes remains an open goal (Sun et al., 2025b), the weight of evidence supports a constructive view: with sufficient optimization depth, bias-aware rewards, and the right domain coverage, RLVR can broaden reasoning capabilities.
Reporting note.
Answer-only metrics (e.g., pass@k) can show muted gains, whereas process-aware criteria such as CoT-pass@k—which require both a correct answer and a syntactically valid chain—often reveal larger deltas in settings explicitly optimized for process rewards (Wen et al., 2025b). We treat these as behavioral signals rather than explanations: chains need not be faithful to internal computation (Chen et al., 2025d). Accordingly, we report both answer- and process-aware metrics alongside calibration (ECE/abstention), and we prefer budget-matched saturation curves over single-point pass@k (Section 4).
Cross-Lens Lessons:
Across both perspectives, two themes are consistent: (i) the base model’s strength and coverage strongly condition observed gains, and (ii) reward/metric design and evaluation protocol (sampling budgets, seeds, and contamination hygiene) can change the sign of conclusions. These observations motivate our stance: treat reliability, safety, and measurement as first-class objectives. The following sections detail the RLVR tax (Section 3), show how evaluation choices inflate or erode claims (Section 4), and present a unified, tax-aware training and evaluation protocol that yields more trustworthy estimates and more transferable improvements (Section 6).
3 Perspective I — The RLVR Tax: Benefits, Costs, and Controls
RLVR reliably lifts accuracy on verifiable tasks, but those gains are easy to misread if we ignore systematic costs—the RLVR tax. We focus on three recurrent pressures: (i) hallucination and overconfidence, (ii) erosion of instruction following, and (iii) safety and privacy exposure. Each maps directly to our broader thesis: without tax-aware training and sober measurement, reported “reasoning gains” are easy to overstate and hard to transfer.
Table 1: Factual-QA control tracking Abstention ($\downarrow$), Shared Accuracy ($\uparrow$), and ECE ($\downarrow$). Interpretation. “Not attempted” is the absolute number of items with no extractable answer, computed on the block-specific evaluation set for that family (totals differ across families). At 14B/32B, reasoning SFT sharply reduces abstentions but leaves shared accuracy flat and worsens calibration; at a larger scale (DeepSeek), RL reduces abstention and improves shared accuracy and ECE, largely by attempting more items. Accuracy on newly attempted tail items is modest.
Family/Scale Model (version) Not attempted $\downarrow$ Accuracy (shared, %) $\uparrow$ ECE (shared) $\downarrow$
Qwen2.5 14B Qwen2.5-14B-Instruct 1136 12.5 0.598
R1-Distill-Qwen-14B (SFT) 102 10.5 0.692
RL-Reasoning-14B 103 10.0 0.684
Qwen2.5 32B Qwen2.5-32B-Instruct 2492 17.4 0.591
R1-Distill-Qwen-32B (SFT) 76 17.5 0.640
RL-Reasoning-32B 63 17.1 0.600
DeepSeek DeepSeek-V3 (Instruct) 480 27.5 0.496
DeepSeek-R1 (RL) 81 34.6 0.317
3.1 Hallucination and Overconfidence
RLVR can suppress refusals while amplifying confident errors. Empirically, refusal rates often collapse after RLVR, shifting abstentions into assertive answers (Song et al., 2025); models may also repeat flawed steps or produce CoT that diverges from the final answer even when accuracy rises (Yao et al., 2025). A mechanistic reading is that sparse, verifiable rewards combined with entropy pressure encourage determinacy under weak evidence, producing high-variance gradients and spurious local optima (Li & Ng, 2025). Consistent with this, multiple studies find substantial miscalibration: self-reported confidence remains high even on incorrect responses and can increase with longer CoT traces (Mei et al., 2025; Zeng et al., 2025; Kirichenko et al., 2025).
We make this concrete with a factual-QA control that tracks Not attempted (lower is better), Accuracy (shared)—accuracy on items that both models at a given scale attempted (higher is better)—and Expected Calibration Error (ECE; lower is better) computed on shared items from the model’s stated confidence (10 uniform bins; see section 6 (C) for details). The dominant effect is attempt inflation: stronger RLVR policies attempt far more questions. For example, moving from DeepSeek-V3 to DeepSeek-R1 (Guo et al., 2025) sharply reduces abstentions and improves shared ECE, but accuracy on the newly attempted tail is modest, so reported gains depend on how one aggregates across attempted vs. unattempted items (Zeng et al., 2025). In other words, scaling RL encourages attempts; whether this yields net quality improvements depends on calibration and gating.
In practice, we mitigate these effects by (i) decomposing reward into correctness, grounding/citation sufficiency, and calibrated refusal (see section 6 (A)); (ii) adding calibration gates (ECE/entropy) and early stopping when confidence outpaces accuracy (see section 6 (C)); and (iii) reporting ECE and refusal alongside accuracy, plus shared-accuracy and attempt-aware summaries (see section 6 (D)).
3.2 Instruction Following
Optimizing hard for reasoning can erode controllability—especially for long generations. Several studies observe regressions in instruction fidelity when training emphasizes extended chain-of-thought or purely verifiable endpoints (Fu et al., 2025; Li et al., 2025). Moreover, recent evidence shows that models tend to overfit to a small set of verifiable constraints and struggle to generalize; IFBench (58 out-of-domain constraints) documents this gap and finds that RLVR improves precise instruction-following generalization (Pyatkin et al., 2025). We counter this by making format/constraint adherence a first-class term in the reward (see section 6 (A)) and by alternating objectives or schedules so the model periodically re-anchors to instruction faithfulness rather than overfitting to long-CoT trajectories (see section 6 (B)).
3.3 Safety and Privacy
Long, explicit reasoning traces widen the attack and leakage surface. Frontier reasoning models are jailbreakable at high rates; automated attacks approach deterministic success in the studied settings, and success rises with attempt and context budgets (Kassianik & Karbasi, 2025; Zhou et al., 2025; Jiang et al., 2025). Many-shot evaluations show the same scaling. Targeted defenses tuned specifically to that regime can sharply reduce measured attack success in controlled tests, but these gains are brittle across models and budgets and can impose utility costs (Ackerman & Panickssery, 2025). Richer CoT also increase privacy risk: attributes, confidential prompts, and dataset contents can be reconstructed more easily as intermediate reasoning grows (Green et al., 2025). Safety tuning itself can impose a “safety tax,” degrading math and coding unless staged and balanced carefully (Huang et al., 2025; Zhang et al., 2025b). Our practice is to cap best-of-$N$ and chain-of-thought length at evaluation and to filter or redact private content in traces (see section 6 (C)), while reporting many-shot jailbreak stress tests and a clear refusal policy as part of standard evaluation (see section 6 (D)).
Takeaway:
These taxes distort reported gains in predictable ways: overconfident hallucinations inflate apparent utility, instruction drift undermines deployability, and expanded attack surfaces raise real-world risk. The answer is not to abandon RLVR but to co-optimize for reliability—compose rewards so correctness, grounding, and abstention can all improve; manage variance and difficulty; and make calibration and provenance part of evaluation. We follow this recipe in the next sections to show where conclusions change once the tax is priced in (section 4, section 6).
4 Perspective II — Pitfalls in Evaluation: Are We Measuring Real Reasoning Gains Accurately?
Claims about RLVR progress are highly sensitive to sampling budgets, metric design, and dataset hygiene. This matters for our thesis: budget mismatch, fragile judge pipelines, and calibration drift can overstate “reasoning gains” and complicate transfer beyond math and code. We treat contamination separately in Section 5 and focus here on budgeting, metric robustness, and calibration.
4.1 Budget Parity and Saturation
A recurring pattern in the literature is to report pass@k for RLVR models while holding base models to much smaller budgets (e.g., pass@1 or pass@5). In such settings, measured improvements often reflect extra search rather than a better policy. Prior work stresses matched budgets and decontaminated baselines for fair comparison (Wen et al., 2025a; Wu et al., 2025b). Small or underspecified benchmarks inflate estimator variance and reduce statistical power, making single-run point estimates unstable; averaging across multiple randomized trials materially improves estimate reliability (Mu et al., 2025). Under parity controls (e.g., SoberScore, a standardized, matched-budget, multi-seed evaluation that fixes decoding/setup and reports mean ± std), several prominent gaps shrink or disappear (Hochlehnert et al., 2025). In practice, evaluations should match $k$ across base and RLVR, plot saturation curves (accuracy vs. $k$, optionally summarized by area under the curve), and disclose decoding budgets and parameters.
4.2 Metric Fragility and LLM Judges
While RLVR often targets verifiable domains (math/coding), many evaluation targets either lack programmatic verifiers or have only partial/weak ones—e.g., safety/refusal appropriateness, long-form coherence and rationale quality, multi-turn task success, or fuzzy information extraction. In those cases, we fall back to LLM-as-a-judge. LLM-judge pipelines offer convenience but can be brittle: small changes—seed choice, $k$, dataset versioning, instruction placement, option order, or even tensor-parallel settings—produce swings comparable to reported gains (Sun et al., 2025a). Judges are also manipulable (Zhao et al., 2025b). Beyond judges, seemingly modest metric or setup shifts can alter conclusions about “stable reasoning” (Liu et al., 2024a). Where programmatic verifiers exist, they should be preferred. When judges are unavoidable, prompt/format perturbations and adversarial probes should be part of the protocol, with robustness deltas reported, inter-judge agreement documented, and all templates/configurations released to reduce hidden degrees of freedom.
4.3 Calibration Drift and Overconfidence
RL optimization often sharpens the output distribution. Top-1 accuracy can rise even as calibration worsens, increasing brittleness under distribution shift (Hochlehnert et al., 2025). Combined with the volatility noted above (Sun et al., 2025a), some apparent “wins” likely reflect confident exploitation of evaluation quirks rather than robust reasoning. Evaluations should therefore track expected calibration error (ECE), output entropy, and refusal/abstention alongside accuracy, and consider early stopping or annealing when calibration degrades even if reward continues to increase. See section 6(D) for our unified checklist.
Table 2: Reported scores vs. standardized evaluation (avg@32 of pass@1, estimated by averaging 32 independent single-sample decodes) with matched decoding budgets and a shared verifier/prompt family. $\Delta$ denotes Reported - Standardized. See Appendix 4 for comprehensive standardized results of recent RLVR models.
Model (checkpoint) Benchmark Reported (setup) Standardized Eval $\Delta$
AIME-24 48.13 (pass@1; Liu et al. 2025a) 45.62 +2.51+2.51
AIME-25 33.33 (pass@1; Liu et al. 2025a) 33.85 0.52-0.52
Nemotron-Research- AMC-23 79.29 (pass@1; Liu et al. 2025a) 85.70 6.41-6.41
Reasoning-Qwen-1.5B v1 Math 91.89 (pass@1; Liu et al. 2025a) 92.01 0.12-0.12
Minerva 47.98 (pass@1; Liu et al. 2025a) 39.27 +8.71+8.71
Olympiad 60.22 (pass@1; Liu et al. 2025a) 64.56 4.34-4.34
AceReason-Nemotron-14B AIME-24 78.60 (avg@64; Yang et al. 2025) 77.29 +1.31+1.31
AIME-25 67.40 (avg@64; Yang et al. 2025) 66.04 +1.36+1.36
DAPO-Qwen-32B AIME-24 50.00 (accuracy; Yu et al. 2025) 51.56 1.56-1.56
Open-RS3-1.5B AIME-24 46.70 (pass@1; Dang & Ngo 2025) 30.94 +15.76+15.76
STILL-3-1.5B AIME-24 39.33 (avg@64; T=0.6T{=}0.6, top-p=0.95p{=}0.95; Min et al. 2025) 31.46 +7.87+7.87
DeepScaleR-1.5B AIME-24 43.10 (pass@1; Luo et al. 2025) 38.54 +4.56+4.56
Polaris-7B-Preview AIME-24 72.60 (avg@32; NLP & collaborators 2025) 66.46 +6.14+6.14
AMC-23 89.00 (avg@8; NLP & collaborators 2025) 93.59 4.59-4.59
4.4 Reported vs. Reproduced: a Gap Analysis
To make these issues concrete, we compared widely cited checkpoints to parity-controlled runs using the same verifier, matched decoding budgets, and aligned dataset versions. The gaps in Table 2 arise from three main factors:
Sampling budgets. Multi-sample reporting (e.g., avg@64 with T=0.6T{=}0.6, top-p=0.95p{=}0.95 for STILL-3-1.5B) inflates scores relative to single-shot pass@1 under the same verifier (Min et al., 2025). In our standardized runs this manifests as +7.87+7.87 for STILL and +6.14+6.14 for Polaris-7B on AIME-24 (NLP & collaborators, 2025).
Template and version drift. Changes in prompt templates and dataset slices (e.g., AIME curation, AMC answer formats) shift accuracy by several points, with exact hashes frequently missing from model cards (NLP & collaborators, 2025; Liu & contributors, 2025). This helps explain mixed signs for Nemotron-1.5B v1 (Minerva +8.71+8.71 vs. AMC-23 6.41-6.41).
Metric and decision rules. Sources sometimes mix “accuracy,” pass@k, avg@k (mean of pass@1 across draws), and maj@k (majority vote), or substitute judges where verifiers exist. Once we normalize to a single verifier/spec, several deltas compress (e.g., DAPO-32B on AIME-24: 1.56-1.56(Sun et al., 2025a; Zhao et al., 2025b).
Finally, small-set variance matters: on AIME-24/25, seed and decode settings alone can yield $\pm$3–5 percentage points. Confidence intervals and saturation curves are needed to separate noise from effect (Mu et al., 2025; Hochlehnert et al., 2025).
Why this matters
Under unmatched budgets, unprobed judges, and untracked calibration, distributional sharpening can masquerade as capability expansion. Clean, parity-controlled evaluation often shrinks or flips celebrated gains. Hence our checklist—budget parity and saturation curves, variance disclosure, judge robustness, calibration metrics, and a contamination audit (Section 5)—is essential to distinguish real progress from artifacts and to make the tax-aware protocol credible.
5 Perspective III — Data Contamination
Data provenance is a first-order confound: if pretraining or RL data overlap evaluation sets, measured “reasoning” may reflect memorization. Contamination-aware evaluations in coding (Liu et al., 2024b) and multi-domain leaderboards (Liang et al., 2023) already recommend strict provenance checks; our partial-prompt audit provides direct evidence in math.
Table 3: Partial–prompt contamination summary. The Math block (left) reports answer–match accuracy at an 80% prefix on two legacy sets vs. the fresh AIME–2025. The SimpleQA block (right) is a non–math control, where Qwen’s advantage largely attenuates. Full results, including ACC@60/40 and ROUGE–L/EM for all models and datasets, are in Tables 610 in Appendix.
Model Math (ACC@80) SimpleQA
MATH-500 AMC-23 AIME-2025 R@80 EM@80
Llama-3-1.8B 2.8 0.0 0.0 37.11 19.86
Qwen2.5-Math-7B 58.0 52.5 0.0 29.34 12.47
Qwen2.5-32B 60.0 52.5 3.3 41.06 23.09
Qwen3-14B-Base 58.2 47.5 0.0 37.24 19.40
Probe design (partial-prompt completion).
Following Wu et al. (2025b), we reveal only the first $x\%$ of each problem ($x\!\in\!\{80,60,40\}$), greedily decode the suffix, and compute ROUGE-L@x, EM@x, and ACC@x. High EM/ACC under large prefixes indicates the model can reconstruct the hidden tail and final answer verbatim.
Case study: Qwen vs. Llama.
As summarized in Table 3, Qwen2.5–Math–7B and Qwen3–14B–Base achieve high ACC@80 on legacy math sets (around 58–60% on MATH-500) yet collapse on AIME–2025 (0–3%). Llama–3–1.8B remains near zero across the same sets, consistent with clean evaluation. This pattern—strong on older math sets, absent on the newest release—aligns with contamination in Qwen pretraining and clean evaluation for Llama. (See App. Tables 69 for ACC/ROUGE/EM at 80/60/40% prefixes.) Similar “contamination-free” stance and methodology are advocated in code by Liu et al. (2024b) and in broader evaluations by Liang et al. (2023).
SimpleQA control.
On a non–math control (SimpleQA) there is no systematic Qwen advantage: for example, Qwen2.5–32B attains 41.06/23.09 (R@80/EM@80) versus Llama–3–1.8B at 37.11/19.86 (App. Table 10). The attenuation on an unseen domain supports the interpretation that elevated partial–prompt math scores on legacy sets reflect contamination rather than a general suffix–reconstruction ability.
Controls and implications.
We (i) treat contaminated sets as probes of memorization, not reasoning; (ii) prioritize uncontaminated or freshly released test sets; and (iii) publish prompts/seeds/configurations to enable third-party screening. When we re-evaluate RLVR vs. base models under these controls, widely cited gaps shrink or flip. This directly supports our thesis: contamination can make modest distributional sharpening appear as frontier expansion.
6 A Tax-Aware RLVR Protocol
Guided by the RLVR “tax” framing (§3), the evaluation pitfalls analysis (§4), and the contamination audit (§5), we specify a tax-aware protocol that preserves RLVR gains while controlling hidden costs. It is designed to curb failure modes yet remain easy to adopt in current pipelines.
(A) Separate what you care about in the reward.
We use componentized rewards that explicitly score three things: answer correctness, grounding/citation sufficiency, and calibrated refusal when evidence is weak. In practice, models train more stably when these components are scheduled, not blended from the start: learn to answer first, then add grounding, and finally layer in refusal once accuracy plateaus or overconfidence rises (Sim et al., 2025; Ren et al., 2025). If a learned judge is unavoidable, measure its fragility via small prompt or format perturbations and “one-token” hacks and report robustness deltas (Zhao et al., 2025b).
(B) Control variance and match difficulty to capacity.
Concentrate exploration at high-entropy forks early, then anneal (Cheng et al., 2025a; Agarwal et al., 2025; Cui et al., 2025; Wang et al., 2025a). Reduce waste via selective sampling: filter trivial or unsolvable prompts and down-sample near-deterministic rollouts (Xiong et al., 2025; Xu et al., 2025; Zheng et al., 2025). Use light negative reinforcement to suppress spurious high-reward modes (Zhu et al., 2025). Match difficulty to capacity by switching between imitation and GRPO via an online hardness signal (e.g., GHPO) (Liu et al., 2025d). Monitor KL to the base model, sequence entropy, and reward variance, and cool down or early-stop on variance collapse or KL spikes.
(C) Mix offline and online; stop by calibration, not reward alone.
Strong offline rollouts are an efficient bootstrap; we switch to online sampling when the offline set saturates. Crucially, we stop or cool down by calibration rather than reward alone: we track expected calibration error (ECE), output entropy, and refusal rates alongside accuracy, and we anneal or halt when ECE worsens or abstention collapses even if reward continues to rise (Hochlehnert et al., 2025; Leng et al., 2024). For deployment, we cap best-of-$N$ attempts and chain-of-thought length, and we consider “think-when-needed” policies to avoid unnecessary long reasoning traces that expand the attack and privacy surface (Fang et al., 2025).
(D) Evaluation and reporting are part of the protocol, not an afterthought.
This protocol standardizes five practices: (1) match sampling budgets for base and RLVR and publish saturation curves (accuracy vs. $k$) with area-under-curve summaries, avoiding pass@k apples-to-oranges comparisons; (2) disclose variance ($\geq$3 seeds with 95% CIs) and the decoding/TP/DP settings that can move results (Jordan et al., 2020; Patterson et al., 2024); (3) track calibration and abstention (ECE, entropy, refusal policy/rates) alongside accuracy (Hochlehnert et al., 2025; Leng et al., 2024); (4) when an LLM judge is used, report judge robustness under prompt/format perturbations and adversarial sweeps (Zhao et al., 2025b); and (5) audit contamination via fuzzy/lexical matches and partial-prompt reconstruction (ACC/EM/ROUGE at 80/60/40%), including a clean held-out set (e.g., AIME-2025) before claiming reasoning gains (Wu et al., 2025b).
(E) A compact default recipe.
In code, we instantiate the reward as
$\mathcal{R}=\alpha\,\text{Correct}+\beta\,\text{Grounding}+\gamma\,\text{Refusal}$
starting with correctness only, adding grounding midway, and introducing refusal once accuracy stalls or ECE climbs. Exploration begins hot and cools (higher temperature or entropy bonus $\rightarrow$ lower), with token-local entropy targeting at uncertain tokens and down-sampling of near-deterministic rollouts; optional negative sampling is applied when spurious high-reward modes appear (Zhu et al., 2025). Objective selection switches online between imitation and GRPO using a running hardness signal (small-$k$ pass@k or semantic entropy) (Liu et al., 2025d; Chen et al., 2025b). A calibration gate halts or anneals when ECE exceeds the baseline by a fixed margin or refusal falls below a floor—even if reward improves. The evaluation pack includes matched budgets and saturation curves, $\geq$3 seeds with CIs, calibration/entropy/refusal metrics, judge robustness deltas (if used), and a contamination audit with a clean out-of-distribution set (Hochlehnert et al., 2025; Sun et al., 2025a; Wu et al., 2025b).
(F) Where this already works in the wild.
Recent mechanisms illustrate the spirit of this protocol. ProRL uses prolonged RL with KL control and periodic resets to broaden support—effective, but calibration-sensitive (Liu et al., 2025a). He et al. (2025a) add unlikeliness rewards to counter rank bias and escape distribution sharpening. GHPO adapts the objective by difficulty to stabilize training (Liu et al., 2025d). Light-R1 shows that curricula spanning SFT/DPO/RL reduce variance and overclaiming (Wen et al., 2025a). SoberScore enforces budget parity and normalization, shrinking prominent gains (Hochlehnert et al., 2025). ThinkLess trims unnecessary chain-of-thought, reducing safety and privacy exposure with minor accuracy cost (Fang et al., 2025). And contamination audits via partial-prompt probes help distinguish memorization from reasoning (Wu et al., 2025b).
We encourage adopting these choices to re-anchor RLVR to reliability: they keep the practical wins while reducing the degrees of freedom through which budget, metric, and contamination artifacts can masquerade as genuine reasoning gains.
7 Conclusion
RLVR delivers real gains on verifiable tasks, but the field often over-indexes on headline accuracy while under-weighting taxes (hallucination/overconfidence, instruction drift, safety & privacy exposure) and measurement (budget mismatch, judge fragility, calibration drift, contamination). Under parity controls and calibration tracking, several celebrated “reasoning gains” shrink, suggesting that part of the progress reflects distributional sharpening rather than durable expansion.
Broader relevance. The same pressures arise in SFT, RLHF, and test-time compute (e.g., CoT and maj@k). Our protocol—budget parity with saturation curves, calibration/abstention metrics, judge robustness checks, and contamination audits—applies unchanged beyond RLVR.
Why focus on RLVR. We center RLVR because programmatic verifiers make the measurement problem tractable and reveal where gains come from. As a result, it is a useful proving ground for reliability methods that transfer to broader LLM training and deployment. We hope this synthesis helps the community separate genuine capability expansion from artifacts of budgets, metrics, and data, and encourages tax-aware training and reporting across RLVR, SFT, and RLHF.
Ethics Statement
This paper raises some potential ethics concerns under the ICLR Code of Ethics and describes the steps we took to mitigate them:
Privacy & leakage. Long, explicit chains-of-thought can reveal sensitive content or training artifacts. We evaluate only on public math/code benchmarks and do not use private prompts. For any illustrative traces, we sanitize identifiers and redact potentially sensitive strings. Our standardized evaluation caps best-of-$N$ and CoT length to reduce leakage surfaces.
Dual-use: jailbreaks & misuse. Methodological details about jailbreak stress tests could be repurposed to increase harm. We report aggregate metrics and robustness deltas, but do not release attack payloads or automation scripts that would materially increase misuse without corresponding defenses. Our evaluation protocol co-optimizes calibrated abstention alongside accuracy.
Measurement integrity. Selective budgeting, fragile judge pipelines, or metric switching can misrepresent results. To reduce this risk, we use matched decoding budgets, fixed verifier/prompt families, multi-seed averaging, and release configurations (decoding parameters, seeds, templates) sufficient to reproduce our reported numbers.
Reproducibility Statement
We make our results reproducible by fixing verifiers and a shared prompt family (section 4), matching decoding budgets and reporting decoding parameters, and averaging over three seeds (Appendix 4, Table2; Figure 2). Datasets and methodology for contamination screens and judge-robustness probes referenced in section 5 are provided with instructions to re-run the audits. Hardware details and aggregate compute ($128{\times}$H100 (96 GB) cluster and $\approx$3,500 GPU-hours) are included to support cost and throughput replication.
Appendix A Appendix
Conventions and metrics.
pass@k = probability at least one of $k$ samples is correct; avg@k = mean pass@1 across $k$ draws; maj@k = majority vote over $k$; ECE = expected calibration error. Arrows ($\uparrow$/↓) indicate whether higher/lower is better.
Usage of Large Language Models.
We used an LLM assistant as a productivity tool for light editing and formatting: smoothing wording the authors had written, polishing captions, refactoring tables/macros, and drafting small utility snippets (e.g., plotting and CSV loaders) that we then modified and verified. All research ideas, evaluation designs, hyperparameters, and analysis decisions are by the authors, and every experiment and number reported was run and checked by us.
Appendix B Potential Criticisms and Responses
On capability expansion.
RLVR can expand capability under specific regimes (e.g., prolonged RL with resets and KL control; unlikeliness-based shaping; domain-targeted RLVR with sparse pretraining) (Liu et al., 2025a; He et al., 2025a; Cheng et al., 2025b). Our claim is not negation but measurement: without tax-aware training and parity-controlled evaluation, expansion is often overstated. The protocol in §6 is compatible with these positive regimes while curbing overclaiming.
On “it’s just data.”
Data quality and curricula matter (Chen et al., 2025a; Wen et al., 2025a), but reward/metric design independently shapes failure modes (determinacy/overconfidence under judge rewards; terse/off-topic outputs under naive factuality; calibration drift) (Chen et al., 2025c; Leng et al., 2024; Hochlehnert et al., 2025). We treat both data and objectives as first-class knobs (see §6).
On budgeted pass@k.
We use pass@k for practical relevance, but ask for matched budgets and saturation curves. Under these controls, headline gaps often shrink or flip (Hochlehnert et al., 2025). Reporting AUC alongside pass@k keeps comparisons fair (§4).
On safety/privacy risk.
Jailbreak success scales with attempts/context; many-shot tuning can reduce measured rates in controlled settings (Zhou et al., 2025; Jiang et al., 2025; Ackerman & Panickssery, 2025). Longer CoT increases leakage surfaces (Green et al., 2025). Our protocol caps Best-of-NN/CoT length at eval and co-optimizes abstention/privacy with accuracy (§6).
On reliance on LLM judges.
When verifiers are unavailable or partial, we use LLM judges with robustness probes (prompt/format perturbations, adversarial tests) and publish templates/configs; judges are manipulable, so robustness deltas are reported (Zhao et al., 2025b; Sun et al., 2025a). Prefer verifiers when possible (§4).
Scope.
Our synthesis centers on verifiable math/code/QA with open-weight models. Agentic/multimodal settings introduce additional privacy/safety channels and may need tailored verifiers and audits; we expect core principles (budget parity, calibration, abstention, contamination screens) to transfer with domain-specific adaptations.
Appendix C Standardized Evaluations
We report full results under matched budgets, fixed verifiers, and a shared prompt family. These tables extend Section 4 and underlie the saturation-curve argument: once $k$, templates, and dataset versions are controlled, several headline gaps narrow. All results in this section use the same verifier and prompt template family; we specify $k$, number of seeds, and dataset versions below.
Key observations.
[…]
Table 4: Standardized evaluation across math benchmarks (higher is better). Styling: within each size family, the best score for a given benchmark is bold and the second-best is underlined. Context: Parity-controlled scores used in Section 4. Setup: pass@1 estimated by averaging over k=32k{=}32 independent single-sample decodes (avg@32), same verifier and prompt family, 3 seeds. Compute: evaluations ran on a $128{\times}$H100 (96 GB) cluster and consumed $\approx$3,500 GPU-hours.
[…]
Qwen2.5-Math-1.5B 8.33 6.35 44.06 66.67 18.42 30.74 29.10
Qwen2.5-Math-1.5B-Instruct 10.10 8.85 55.08 74.83 29.32 40.00 36.37
DeepSeek-R1-Distill-Qwen-1.5B 31.15 24.06 72.81 85.01 32.18 51.55 49.46
STILL-3-1.5B 31.46 25.00 75.08 86.24 32.77 53.84 50.73
DeepScaleR-1.5B 38.54 30.52 80.86 88.79 36.19 58.95 55.64
Qwen2.5-Math-1.5B-Oat-Zero 20.00 10.00 52.50 74.20 26.84 37.78 36.89
Open-RS1 30.94 22.60 73.05 84.90 29.92 52.82 49.04
Open-RS2 28.96 24.37 73.52 85.06 29.74 52.63 49.05
Open-RS3 30.94 24.79 72.50 84.47 29.11 52.25 49.01
Nemotron-Research-Reasoning-Qwen-1.5B v1 45.62 33.85 85.70 92.01 39.27 64.56 60.17
Nemotron-Research-Reasoning-Qwen-1.5B v2 51.77 32.92 88.83 92.24 39.75 64.69 61.70
Qwen2.5-Math-7B 15.62 6.56 52.81 67.72 15.64 32.44 31.80
Qwen2.5-Math-7B-Instruct 12.19 9.17 58.36 83.21 35.56 41.60 40.01
DeepSeek-R1-Distill-Qwen-7B 53.23 38.96 89.30 93.95 43.07 66.67 64.20
Qwen2.5-Math-7B-Oat-Zero 26.67 6.67 67.50 79.20 32.72 41.78 42.42
Skywork-OR1-7B 66.88 51.15 92.73 96.04 44.03 73.61 70.74
LEAD-7B 51.67 37.19 89.06 93.73 43.11 66.34 63.52
AceReason-Nemotron-7B 65.83 47.19 95.08 95.81 45.35 73.92 70.53
AceReason-Nemotron-1.1-7B 71.56 64.58 93.36 96.73 44.05 76.37 74.44
Polaris-7B-Preview 66.46 51.56 93.59 95.68 44.47 73.65 70.90
Qwen2.5-14B 11.04 7.92 47.19 73.19 22.51 37.01 33.14
Qwen2.5-14B-Instruct 13.65 12.40 58.13 80.28 38.63 43.23 41.05
DeepSeek-R1-Distill-Qwen-14B 67.81 48.33 95.39 95.28 46.43 72.06 70.88
LEAD-14B 64.06 52.29 92.81 95.23 47.52 72.25 70.69
AceReason-Nemotron-14B 77.29 66.04 98.67 96.90 47.73 77.34 77.33
Qwen2.5-32B 15.62 9.17 59.30 76.51 26.42 41.45 38.08
Qwen2.5-32B-Instruct 17.19 14.17 67.66 83.17 40.96 47.85 45.17
DeepSeek-R1-Distill-Qwen-32B 69.06 55.52 95.62 95.74 46.50 72.76 72.53
DAPO-Qwen-32B 51.56 36.98 92.73 80.74 33.07 48.97 57.34
Enigmata-Qwen2.5-32B 61.67 46.88 91.25 93.69 46.32 69.14 68.16
DeepSeek-V3-0324 55.83 43.33 92.50 95.12 48.58 66.91 67.05
DeepSeek-R1-0528 90.00 78.33 99.38 97.80 48.35 76.44 81.72
Qwen3-4B 21.88 17.92 66.95 84.27 38.50 52.38 46.98
Qwen3-4B (Thinking) 72.40 62.60 95.78 96.31 46.44 76.40 74.99
Polaris-4B-Preview 27.29 23.02 71.56 85.89 39.20 59.13 51.02
Polaris-4B-Preview (Thinking) 79.90 77.71 99.45 97.38 47.21 80.38 80.34
Qwen3-8B 25.42 20.21 69.61 84.54 39.81 54.15 48.96
Qwen3-8B (Thinking) 77.29 66.46 95.00 96.86 49.06 77.56 77.04
DeepSeek-R1-0528-Qwen3-8B 75.73 67.29 97.19 96.32 45.85 74.19 76.10
Appendix D Sequence Length and Compute Footprint
We summarize generated token lengths by benchmark as a proxy for compute footprint and exposure surface. Longer chains increase latency and may worsen calibration and safety/privacy risks (Section 3), So we report typical generation lengths under the same decoding budgets used in Appx. C.
Table 5: Median generated output length (tokens per problem; answer + reasoning) under the standardized decoding setup. Context: Compute/privacy footprint complementing accuracy. Setup: Same prompts/verifier as Appx. C; decoding budget k=32k{=}32; lengths aggregated across all samples per problem (correct and incorrect). Takeaway: Reasoning-tuned models often produce much longer traces (e.g., 2–8× vs. instruct baselines), which impacts evaluation cost and safety/privacy exposure.
Model AIME-24 AIME-25 AMC-23 Math Minerva Olympiad Avg
Qwen2.5-Math-1.5B 1114 1033 828 648 959 904 914
Qwen2.5-Math-1.5B-Instruct 990 891 801 566 640 810 783
DeepSeek-R1-Distill-Qwen-1.5B 16363 16252 9979 5700 8194 11873 11394
STILL-3-1.5B 13350 13000 7716 4314 5921 9345 8941
DeepScaleR-1.5B 9780 8978 5003 3139 5270 5807 6330
Qwen2.5-Math-1.5B-Oat-Zero 1166 1155 848 626 689 892 896
Open-RS1 13578 13698 7495 4170 5935 8995 8978
Open-RS2 14215 13618 7612 4174 5805 9059 9081
Open-RS3 14394 13606 7884 4233 5688 9047 9142
Nemotron-Research-Reasoning-Qwen-1.5B 7786 7713 6294 5070 6569 6678 6685
Qwen2.5-Math-7B 1201 1152 946 720 1136 950 1018
Qwen2.5-Math-7B-Instruct 1465 1418 983 670 748 1051 1056
DeepSeek-R1-Distill-Qwen-7B 13613 14543 6402 4125 5595 8988 8878
Qwen2.5-Math-7B-Oat-Zero 1032 1178 884 673 696 870 889
Skywork-OR1-7B 15366 17845 8361 5541 8566 11818 11250
LEAD-7B 10838 11573 4863 3111 3692 7054 6855
AceReason-Nemotron-1.1-7B 14331 16502 6672 3835 6676 9060 9513
Polaris-7B-Preview 12564 14389 6538 4313 6125 8681 8768
Qwen2.5-14B 1076 1088 847 591 1209 815 938
Qwen2.5-14B-Instruct 1079 994 850 607 632 844 834
DeepSeek-R1-Distill-Qwen-14B 11295 13389 5735 3781 4919 8042 7860
LEAD-14B 8364 9019 4469 3031 4675 5713 5879
AceReason-Nemotron-14B 13871 16334 7239 4609 7677 10030 9960
Qwen2.5-32B 1165 1055 782 556 1186 800 924
DeepSeek-R1-Distill-Qwen-32B 10979 13012 5826 3652 4663 7924 7676
DAPO-Qwen-32B 6627 6074 3086 2812 3887 5122 4601
Enigmata-Qwen2.5-32B 13204 15417 10195 6358 13770 12098 11840
DeepSeek-V3-0324 3601 3474 2134 1281 705 2386 2264
DeepSeek-R1-0528 16276 18416 10285 6041 7085 13750 11976
MiMo-7B-Base 21716 18842 15734 6390 13809 13299 14965
MiMo-7B-SFT 13012 14250 7030 4177 6152 8421 8841
MiMo-7B-RL-Zero 18105 18575 9950 6564 6546 12152 11982
Qwen3-4B 8909 6079 2369 1289 812 3217 3779
Qwen3-4B (Thinking) 14750 17813 8290 5076 6682 10606 10536
Polaris-4B-Preview 5582 6148 2637 1293 804 4632 3516
Polaris-4B-Preview (Thinking) 28369 33109 15125 9333 13307 21832 20179
Qwen3-8B 8375 6151 2863 1479 662 3624 3859
Qwen3-8B (Thinking) 14764 18296 8756 5391 7105 11384 10949
DeepSeek-R1-0528-Qwen3-8B 20970 22742 12895 7640 9329 16108 14947
Appendix E Data Contamination Probes: Math
We use partial-prompt completion to detect tail reconstruction: reveal the first $x\%$ of the problem ($x\!\in\!\{80,60,40\}$), greedily decode the remainder, and score ACC/EM/ROUGE-L at that prefix. High scores at large prefixes indicate memorization of legacy sets rather than reasoning. Tables here expand Section 5.
Table 6: Accuracy of Qwen3 checkpoints on five mathematics benchmarks when each model receives only the first $x\%$ of the question (x=80,60,40x\!=\!80,60,40) and must greedily complete the remainder. Columns “ACC (X%)” report the average accuracy at that prefix length. Takeaway: Qwen3 variants achieve high ACC@80 on legacy MATH-500/AMC-23 but collapse on AIME-2025, consistent with contamination in older sets.
Model Dataset ACC (80%) ACC (60%) ACC (40%)
MATH‑500 26.40 14.40 5.40
AMC‑23 20.00 5.00 0.00
Qwen3‑0.6B‑Base AIME 2024 3.33 0.00 0.00
AIME 2025 0.00 0.00 0.00
Minerva‑Math 1.10 0.00 0.00
MATH‑500 33.80 18.40 9.00
AMC‑23 20.00 15.00 0.00
Qwen3‑1.7B‑Base AIME 2024 0.00 0.00 0.00
AIME 2025 3.33 3.33 0.00
Minerva‑Math 2.94 2.21 0.74
MATH‑500 43.20 28.20 15.00
AMC‑23 37.50 37.50 17.50
Qwen3-4B‑Base AIME 2024 10.00 6.67 10.00
AIME 2025 3.33 0.00 0.00
Minerva‑Math 4.78 2.31 0.74
MATH‑500 52.00 35.60 24.80
AMC‑23 35.00 25.00 27.50
Qwen3‑8B‑Base AIME 2024 13.33 6.67 10.00
AIME 2025 3.33 3.33 0.00
Minerva‑Math 5.88 3.31 1.84
MATH‑500 58.20 44.80 30.20
AMC‑23 47.50 37.50 32.50
Qwen3‑14B‑Base AIME 2024 10.00 16.67 16.67
AIME 2025 0.00 0.00 0.00
Minerva‑Math 5.88 2.21 2.94
MATH‑500 62.40 51.40 37.80
AMC‑23 40.00 25.00 25.00
Qwen3‑30B‑A3B AIME 2024 13.33 23.33 13.33
AIME 2025 3.33 0.00 0.00
Minerva‑Math 4.78 3.21 1.84
Table 7: Accuracy of Qwen2.5 and Llama‑3 checkpoints on five mathematics benchmarks when each model receives only the first $x\%$ of the question (x=80,60,40x\!=\!80,60,40) and must greedily complete the remainder. Takeaway: Qwen2.5 shows strong tail reconstruction on legacy math; Llama-3-1.8B remains near zero across sets, reinforcing the contamination interpretation.
Model Dataset ACC (80%) ACC (60%) ACC (40%)
MATH‑500 58.00 44.40 29.20
AMC‑23 52.50 42.50 32.50
Qwen2.5‑Math‑7B AIME 2024 16.67 20.00 16.67
AIME 2025 0.00 0.00 0.00
Minerva‑Math 7.72 4.41 2.94
MATH‑500 44.80 27.80 15.00
AMC‑23 27.50 27.50 22.50
Qwen2.5‑7B AIME 2024 6.67 0.00 3.33
AIME 2025 3.33 3.33 0.00
Minerva‑Math 6.62 4.78 1.47
MATH‑500 44.80 28.60 14.00
AMC‑23 40.00 15.00 12.50
Qwen2.5‑7B‑Instruct AIME 2024 3.33 0.00 0.00
AIME 2025 3.33 0.00 0.00
Minerva‑Math 10.29 5.51 3.31
MATH‑500 50.60 35.80 21.00
AMC‑23 40.00 27.50 27.50
Qwen2.5‑14B AIME 2024 10.00 3.33 6.67
AIME 2025 3.33 0.00 0.00
Minerva‑Math 8.46 6.25 2.21
MATH‑500 60.00 47.80 32.00
AMC‑23 52.50 45.00 42.50
Qwen2.5‑32B AIME 2024 16.67 13.33 10.00
AIME 2025 3.33 0.00 6.67
Minerva‑Math 9.93 4.41 2.94
MATH‑500 2.80 2.40 2.00
AMC‑23 0.00 0.00 0.00
Llama‑3‑1.8B AIME 2024 0.00 0.00 0.00
AIME 2025 0.00 0.00 0.00
Minerva‑Math 1.84 0.37 0.00
Table 8: RougeL and exact‑match (EM) scores for Qwen3.0 checkpoints on five mathematics benchmarks when each model receives only the first $x\%$ of the question (x=80,60,40x\!=\!80,60,40) and must greedily complete the remainder. Columns “R@xx” and “EM@xx” report the average ROUGE‑L and EM for that prefix length. Takeaway: Elevated ROUGE-L/EM at large prefixes on legacy sets (but not on AIME-2025) indicates suffix reconstruction rather than emergent reasoning.
Model Dataset R@80 EM@80 R@60 EM@60 R@40 EM@40
MATH-500 53.31 20.80 45.90 6.40 39.33 1.00
AMC-23 52.26 7.50 40.39 0.00 33.92 0.00
Qwen3-0.6B-Base AIME 2024 54.85 20.00 27.79 0.00 24.76 0.00
AIME 2025 54.69 10.00 33.47 0.00 29.16 0.00
Minerva-Math 31.41 1.47 29.17 0.00 24.86 0.00
MATH-500 55.86 24.60 49.05 10.20 41.91 2.40
AMC-23 64.35 30.00 52.57 20.00 44.95 12.50
Qwen3-1.7B-Base AIME 2024 56.55 26.67 36.48 6.67 36.56 6.67
AIME 2025 53.90 16.67 39.40 6.00 31.09 0.00
Minerva-Math 33.99 3.31 30.32 0.00 27.28 0.00
MATH-500 65.04 37.40 55.98 18.80 46.44 7.00
AMC-23 73.24 52.50 68.56 45.00 63.89 35.00
Qwen3-4B-Base AIME 2024 68.98 43.33 53.22 30.00 53.22 26.67
AIME 2025 55.34 10.00 39.29 0.00 27.72 0.00
Minerva-Math 34.13 2.94 32.19 0.37 28.18 0.00
MATH-500 70.98 47.20 63.69 29.60 53.84 15.20
AMC-23 78.23 57.50 69.09 52.50 69.73 47.50
Qwen3-8B-Base AIME 2024 77.28 60.00 55.76 33.33 57.85 30.00
AIME 2025 51.97 10.00 38.60 0.00 32.21 0.00
Minerva-Math 35.87 2.57 33.41 0.00 28.52 0.00
MATH-500 74.93 55.80 70.08 39.80 60.23 23.80
AMC-23 75.32 55.00 73.20 60.00 75.52 52.50
Qwen3-14B-Base AIME 2024 72.31 50.00 61.48 40.00 60.54 40.00
AIME 2025 56.07 10.00 36.41 0.00 32.46 0.00
Minerva-Math 38.12 4.41 34.11 0.74 29.72 0.00
MATH-500 80.12 62.40 75.03 45.60 64.93 33.00
AMC-23 80.40 60.00 73.24 55.00 77.21 55.00
Qwen3-30B-A3B AIME 2024 74.26 53.33 66.53 40.00 62.90 33.33
AIME 2025 51.57 6.67 43.18 0.00 30.55 0.00
Minerva-Math 37.01 3.68 34.96 1.10 29.22 0.00
Table 9: RougeL and exact‑match (EM) scores for Qwen2.5 and Llama‑3 on five mathematics benchmarks when each model receives only the first $x\%$ of the question (x=80,60,40x\!=\!80,60,40) and must greedily complete the remainder. Takeaway: Qwen2.5 shows strong reconstruction on legacy math; Llama-3 remains low.
Model Dataset R@80 EM@80 R@60 EM@60 R@40 EM@40
MATH‑500 79.63 61.60 70.30 41.20 60.81 25.40
AMC‑23 77.35 57.50 71.47 50.00 66.29 42.50
Qwen2.5‑Math‑7B AIME 2024 69.35 53.33 55.90 30.00 56.53 30.00
AIME 2025 48.20 10.00 36.31 0.00 27.23 0.00
Minerva‑Math 33.42 2.94 30.41 0.37 26.88 0.00
MATH‑500 66.09 38.20 58.15 18.00 48.34 6.00
AMC‑23 69.47 47.50 67.70 45.00 61.08 35.00
Qwen2.5‑7B AIME 2024 61.28 30.00 46.19 16.67 46.14 16.67
AIME 2025 54.62 10.00 42.41 0.00 30.56 0.00
Minerva‑Math 33.88 2.94 31.13 0.00 27.06 0.00
MATH‑500 60.05 28.60 51.10 10.40 43.41 2.60
AMC‑23 61.38 30.00 51.22 17.50 42.90 10.00
Qwen2.5‑7B‑Instr AIME 2024 55.41 20.00 36.28 0.00 30.64 0.00
AIME 2025 58.12 10.00 37.96 0.00 29.81 0.00
Minerva‑Math 34.37 2.94 28.50 0.00 25.99 0.00
MATH‑500 69.50 45.00 61.20 27.00 51.24 12.80
AMC‑23 74.58 47.50 68.20 42.50 67.82 37.50
Qwen2.5‑14B AIME 2024 65.42 40.00 49.76 23.33 48.97 23.33
AIME 2025 52.56 6.67 38.02 0.00 32.92 0.00
Minerva‑Math 35.64 3.68 31.30 0.00 27.93 0.00
MATH‑500 76.98 56.60 71.14 41.20 58.32 25.20
AMC‑23 79.18 62.50 74.59 60.00 75.78 55.00
Qwen2.5‑32B AIME 2024 73.25 56.67 60.92 40.00 60.47 33.33
AIME 2025 53.67 6.67 40.91 0.00 31.90 0.00
Minerva‑Math 34.88 2.94 32.92 0.00 28.32 0.00
MATH‑500 46.49 14.00 38.75 3.00 32.19 0.60
AMC‑23 40.46 0.00 29.87 0.00 25.28 0.00
Llama‑3‑1.8B AIME 2024 49.70 16.67 31.45 0.00 27.13 0.00
AIME 2025 50.16 6.67 32.97 0.00 27.18 0.00
Minerva‑Math 35.32 1.47 29.17 0.00 26.95 0.00
Appendix F Data Contamination Probes: SimpleQA (Control)
To test whether partial-prompt reconstruction reflects a general suffix-completion ability, we repeat the probe on SimpleQA. If math-set effects were general, we would expect similarly high ROUGE-L/EM at large prefixes here. They are not: scores cluster modestly across families.
Table 10: Rouge‑L and exact‑match (EM) scores on SimpleQA when each model receives only the first $x\%$ of the question (x=80,60,40x\!=\!80,60,40) and must greedily complete the remainder. Takeaway: No consistent Qwen advantage; effects seen on legacy math do not generalize, supporting the contamination interpretation.
Model Dataset R@80 EM@80 R@60 EM@60 R@40 EM@40
Qwen2.5‑Math‑7B SimpleQA 29.34 12.47 16.19 0.69 13.28 0.00
Qwen2.5‑7B SimpleQA 37.61 19.63 19.88 1.85 14.64 0.23
Qwen2.5‑7B‑Instruct SimpleQA 36.06 17.78 19.76 1.62 15.87 0.46
Qwen2.5‑14B SimpleQA 39.97 21.94 21.69 3.23 15.98 0.23
Qwen2.5‑32B SimpleQA 41.06 23.09 21.87 3.23 14.78 0.00
Llama‑3‑1.8B SimpleQA 37.11 19.86 20.24 2.08 14.22 0.00
Qwen3‑0.6B‑Base SimpleQA 26.45 10.39 16.35 0.46 14.42 0.00
Qwen3‑1.7B‑Base SimpleQA 27.55 12.47 16.50 0.92 14.97 0.23
Qwen3‑4B‑Base SimpleQA 33.52 16.40 17.87 1.85 14.93 0.00
Qwen3‑8B‑Base SimpleQA 33.99 16.40 19.23 1.39 15.39 0.00
Qwen3‑14B‑Base SimpleQA 37.24 19.40 21.44 2.31 16.65 0.23
Qwen3‑30B‑A3B‑Base SimpleQA 38.80 19.63 22.07 3.00 16.55 0.00
Qwen3‑23B‑A22B‑Instr‑2507 SimpleQA 53.88 38.34 34.58 10.39 22.74 1.39