### v2

# The Hidden Costs and Measurement Gaps of Reinforcement Learning with Verifiable Rewards

Abstract
Reinforcement learning with verifiable rewards (RLVR) is a practical, scalable way to improve large language models on math, code, and other structured tasks. However, we argue that many headline RLVR gains are not yet well validated because reports often conflate policy improvement with three confounds: (i) budget mismatch between RLVR and baseline evaluation, (ii) attempt inflation and calibration drift that convert abstentions into confident answers, and (iii) data contamination in benchmarks. Using budget-matched reproductions and partial-prompt contamination probes, we find that several widely cited gaps shrink substantially or disappear once budgets, prompts, and dataset versions are matched, and contaminated sets are treated as memorization probes rather than evidence of reasoning. This does not mean that RLVR is ineffective, but it implies that current measurements often overstate capability gains and obscure reliability costs. We therefore propose a compact, tax-aware minimum standard for RLVR training and evaluation: budget-matched saturation curves with variance, calibration, and abstention tracking, one judge robustness stress test when LLM judges are used, and an explicit contamination screen. With these controls, RLVR remains effective and deployable in verifiable domains, but reasoning gains should be treated as provisional without them.
The Hidden Costs and Measurement Gaps of Reinforcement Learning with Verifiable Rewards
Fang Wu1,*†   Aaron Tu2,*   Weihao Xuan3,*   Heli Qi4,*   Xu Huang5   Qingcheng Zeng6 Shayan Talaei1   Yijia Xiao7   Peng Xia8   Xiangru Tang9   Yuchen Zhuang5   Bing Hu10 Hanqun Cao11   Wenqi Shi12   Rui Yang14   Nan Liu14   Huaxiu Yao8   Ge Liu13 Li Erran Li15   Amin Saberi1   Naoto Yokoya3   Jure Leskovec1   Yejin Choi1,† 1Stanford University  2UC Berkeley  3The University of Tokyo  4Waseda University 5Georgia Tech  6Northwestern University  7UCLA  8UNC Chapel Hill  9Yale University 10Independent Researcher  11CUHK  12UT Southwestern Medical Center  13UIUC 14National University of Singapore  15Amazon AWS AI *Equal Contribution   Corresponding Authors
1 Introduction
Reinforcement learning with verifiable rewards (RLVR) has become a leading post-training route for improving large language models on math and code tasks (Luong et al., 2024; Wen et al., 2025a). By optimizing against automatically computable signals such as unit tests for programs, exact numeric or string matches for math, or retrieval-grounded checks for citations, RLVR promises a path to better performance. Recent work reports large benchmark gains across domains. Appendix Figure 2 shows a trend plot of RLVR-tagged paper activity and AIME-24/25 performance with striking monthly gains. Our position is that RLVR is effective and deployable for verifiable domains, but “reasoning” gains should be treated as provisional without a small core of tax-aware controls.
A central question that frames the current debate is: Does RLVR genuinely impart new reasoning capability, or does it mainly sharpen selection among behaviors the base model already knows how to produce? Budget-parity controlled studies show that base models can narrow or erase RLVR gaps when given matched sampling budgets, consistent with smarter search rather than capability expansion (Yue et al., 2025; Wu et al., 2025c). At the same time, several regimes report gains that are hard to recover by sampling alone: explicitly optimizing the multi-sample objective (pass@k training) (Wang et al., 2025), curriculum via self-play with variational problem synthesis (Zhang et al., 2025a), and distribution-aware reward shaping that counters rank bias and diversity collapse (Liu et al., 2025b), alongside longer-horizon RL schedules and unlikeliness rewards (Liu et al., 2025a; He et al., 2025a). Metric choice can flip conclusions: answer-only scores may diverge from process-aware metrics that require both a correct answer and a valid chain of thought.
A second complication is what we call the RLVR tax: unintended empirical side effects that ride along with apparent gains under current reasoning-style post-training (reduced abstention, miscalibration, instruction-fidelity drift, and a larger safety/privacy surface due to longer traces). This tax is not inevitable and not unique to RLVR; similar patterns arise under reasoning-heavy SFT and RLHF (Wu et al., 2025b). We focus on RLVR because verifiable objectives and open weights make these trade-offs measurable at relatively low cost. In practice, RLVR tends to reduce abstention and increase stated confidence, sometimes even when answers are wrong, thereby shifting risk from “I don’t know” to assertive errors (Song et al., 2025; Mei et al., 2025; Yao et al., 2025). It can also chip away at instruction fidelity on longer generations, where adhering to formats or constraints becomes harder as chains grow (Fu et al., 2025; Li et al., 2025). Finally, longer, more explicit reasoning traces expand the attack and leakage surface, raising jailbreak success and privacy exposure if left unchecked (Zhou et al., 2025a; Jiang et al., 2025; Ackerman and Panickssery, 2025; Green et al., 2025; Huang et al., 2025; Zhang et al., 2025b).
A third complication is measurement. Reported advances are sensitive to sampling budgets and decoding settings, to the stability of LLM-as-a-judge evaluators, to calibration drift, and to data provenance (Hochlehnert et al., 2025; Chandak et al., 2025; Brown and others, 2024; Muennighoff et al., 2025; Zhao et al., 2025; Wang et al., 2024a, b; Leng et al., 2024; Shen et al., 2025; Ye et al., 2026; He et al., 2025b; Mirzadeh et al., 2024; Liu et al., 2024b; Xu et al., 2025; Wu and others, 2024; Wu et al., 2025a, ). When budgets are matched, judges are stress-tested, and datasets are versioned and decontaminated, reported improvement gaps can diminish, suggesting that part of the apparent progress reflects evaluation design rather than durable capability (Hochlehnert et al., 2025; Chandak et al., 2025). We organize the paper around three threads: the RLVR tax, evaluation pitfalls, and data contamination, culminating in a compact tax-aware minimum standard. A visual roadmap is provided in Figure 1.
Figure 1: Paper Roadmap: taxes, evaluation pitfalls, contamination, and the unified protocol.
Scope and definitions.
We use RLVR to mean post-training that optimizes LLMs against automatically checkable signals rather than human preference models (Ouyang and others, 2022). Typical implementations adopt PPO/GRPO/DAPO-style updates with a KL penalty to the base policy, often mixing offline rollouts with online sampling, adding entropy regularization, and using selective filtering or gating. Rewards are frequently componentized (for correctness, grounding/citation sufficiency, and calibrated refusal) and introduced in stages. Our analysis centers on open-weight models fine-tuned for math/code/question answering (QA) with verifiable objectives, in both single- and multi-domain regimes. We deliberately avoid a catalogue of low-level algorithmic variants except where they bear on reliability (e.g., unlikeliness rewards that counter rank bias, or curricula that change variance). The primary threats considered are privacy leakage from long chains of thought (CoT) and increased jailbreak susceptibility during or after RLVR.
Contributions and Roadmap.
Our contributions are threefold: (1) Measurement: we provide budget-matched reproductions and a unified gap table that isolates from budget mismatch, judge and template drift, and dataset versioning (Table 2). (2) Tax: we quantify attempt inflation and the resulting calibration and refusal costs under RLVR and matched SFT controls (Table 1). (3) Protocol: based on these findings, we distill a compact tax-aware minimum standard for RLVR training and evaluation, intended as a minimum credible reporting bar, instantiated in section 6. The paper is organized accordingly: we (i) synthesize evidence on sharpening vs. expansion (section 2); (ii) analyze the RLVR tax and how it distorts reported improvements (section 3); (iii) reproduce representative gaps under parity controls (section 4); (iv) audit contamination with partial-prompt probes (section 5); and (v) present the tax-aware minimum standard (section 6).
2 What Changes Under RLVR? An Evidence Overview
We summarize evidence under two lenses: sharpening vs broadening. Skeptical and optimistic perspectives recur in the literature, and each tends to appear under identifiable conditions.
2.1 Skeptical Lens: Sharpening and Sampling Effects
RLVR often appears to improve sample efficiency and steer the model toward high-reward regions already present in the base distribution, rather than broaden fundamental capability (Gandhi et al., 2025; Shah and others, 2025). When base models are evaluated under matched sampling budgets (i.e., pass@k), gaps between base and RLVR-trained models often shrink (Yue et al., 2025), consistent with sharpening rather than expansion. Empirically, outputs can collapse toward a dominant pretraining mode with reduced diversity (Zhao and others, 2025), and theory suggests that after RLVR, support shrinkage tends to outweigh support expansion (Wu et al., 2025c). Stress tests add caution: performance collapses at higher complexities have been reported even with ample compute (Shojaee et al., 2025), although some failures trace to evaluation artifacts such as token limits and unsolvable items, underscoring the need for budget parity, seed control, and dataset hygiene (see section 4(Opus and Lawsen, 2025). Finally, several works separate knowledge injection from policy optimization: pipelines that use SFT or distillation to inject knowledge and RLVR to select among candidates typically beat pure GRPO, suggesting RLVR optimizes within learned support (Ma et al., 2025; Wan et al., 2025; Chen et al., 2025a); guidance and self-distillation further help over vanilla GRPO (Nath et al., 2025; Liu et al., 2025c).
2.2 Optimistic Lens: Signals of Capability Broadening
Carefully designed RLVR appears to expand what models can do, not merely sharpen selection. Prolonged optimization with explicit KL control and periodic resets (ProRL) reports trajectories that are not reachable by sampling the base policy alone, with the largest gains where the base initially struggles (Liu et al., 2025a). Reward shaping also matters: by diagnosing a GRPO tendency to reinforce high-probability completions, He et al. (2025a) shows that adding unlikeliness rewards and training deeper widens the model’s support and improves multi-sample theorem proving, indicating qualitative changes to reachable reasoning modes. These effects are conditioned by pretraining coverage: cross-domain analyses find broader benefits in Math/Code/Science, while under-represented areas such as Logic/Simulation/Tabular realize meaningful gains when RLVR is applied in-domain (Cheng et al., 2025). Consistent with this, process-aware metrics that require both a correct answer and a syntactically valid chain (CoT-pass@k) often reveal larger RLVR deltas than answer-only measures (Wen et al., 2025b). While generalization to substantially harder regimes remains an open goal (Sun et al., 2025b), the weight of evidence supports a constructive view: with sufficient optimization depth, bias-aware rewards, and the right domain coverage, RLVR can broaden reasoning capabilities.
Reporting note.
Answer-only metrics (e.g., pass@k) can show muted gains, whereas process-aware criteria such as CoT-pass@k (which require both a correct answer and a syntactically valid chain) often reveal larger deltas in settings explicitly optimized for process rewards (Wen et al., 2025b). We treat these as behavioral signals rather than explanations: chains need not be faithful to internal computation (Chen et al., 2025c). Accordingly, we report both answer- and process-aware metrics alongside calibration (ECE/abstention), and we prefer budget-matched saturation curves over single-point pass@k (section 4).
Cross-Lens Lessons:
Across both perspectives, two themes are consistent: (i) the base model’s strength and coverage strongly condition observed gains, and (ii) reward/metric design and evaluation protocol (sampling budgets, seeds, and contamination hygiene) can change the sign of conclusions. These observations motivate our position.
3 Perspective I: The RLVR Tax: Benefits, Costs, and Controls
RLVR reliably lifts accuracy on verifiable tasks, but those gains are easy to misread if we ignore systematic costs, the RLVR tax. We elaborate on three recurrent pressures: (i) hallucination and overconfidence, (ii) erosion of instruction following, and (iii) safety and privacy exposure. Each maps directly to our position: without tax-aware controls and sober measurement, reported “reasoning gains” are easy to overstate and hard to transfer.
Table 1: Factual-QA control tracking Abstention ($\downarrow$), Shared Accuracy ($\uparrow$), and ECE ($\downarrow$). Interpretation. “Not attempted” is the absolute number of items with no extractable answer, computed on the block-specific evaluation set for that family (totals differ across families). At 14B/32B, reasoning SFT sharply reduces abstentions but leaves shared accuracy flat and worsens calibration; at a larger scale (DeepSeek), RL reduces abstention and improves shared accuracy and ECE, largely by attempting more items. Accuracy on newly attempted tail items is modest.
Family/Scale Model (version) Not attempted $\downarrow$ Accuracy (shared, %) $\uparrow$ ECE (shared) $\downarrow$
Qwen2.5 14B Qwen2.5-14B-Instruct 1136 12.5 0.598
R1-Distill-Qwen-14B (SFT) 102 10.5 0.692
RL-Reasoning-14B 103 10.0 0.684
Qwen2.5 32B Qwen2.5-32B-Instruct 2492 17.4 0.591
R1-Distill-Qwen-32B (SFT) 76 17.5 0.640
RL-Reasoning-32B 63 17.1 0.600
DeepSeek DeepSeek-V3 (Instruct) 480 27.5 0.496
DeepSeek-R1 (RL) 81 34.6 0.317
3.1 Hallucination and Overconfidence
RLVR can suppress refusals while amplifying confident errors. Empirically, refusal rates often collapse after RLVR, shifting abstentions into assertive answers (Song et al., 2025); models may also repeat flawed steps or produce CoT that diverges from the final answer even when accuracy rises (Yao et al., 2025). One hypothesis is that sparse, verifiable rewards combined with entropy pressure tend to encourage determinacy under weak evidence, producing high-variance gradients and spurious local optima (Li and Ng, 2025). Consistent with this, multiple studies find substantial miscalibration: self-reported confidence remains high even on incorrect responses and can increase with longer CoT traces (Mei et al., 2025; Zeng et al., 2025; Kirichenko et al., 2025).
We make this concrete with a factual-QA control that tracks Not attempted (lower is better), Accuracy (shared) (accuracy on items that both models at a given scale attempted, higher is better), and Expected Calibration Error (ECE; lower is better) computed on shared items from the model’s stated confidence. The dominant effect is attempt inflation: stronger RLVR policies attempt far more questions. For example, moving from DeepSeek-V3 to DeepSeek-R1 (Guo et al., 2025) sharply reduces abstentions and improves shared ECE, but accuracy on the newly attempted tail is modest, so reported gains depend on how one aggregates across attempted vs. unattempted items (Zeng et al., 2025). In other words, scaling RL encourages attempts; whether this yields net quality improvements depends on calibration and gating.
3.2 Instruction Following
Reasoning-optimized training can erode controllability, especially for long generations. Several studies observe regressions in instruction fidelity when training emphasizes extended chain-of-thought or purely verifiable endpoints (Fu et al., 2025; Li et al., 2025). Moreover, recent evidence shows that models tend to overfit to a small set of verifiable constraints and struggle to generalize; IFBench (58 out-of-domain constraints) documents this gap and finds that RLVR improves precise instruction-following generalization (Pyatkin et al., 2025). This is why our minimum standard in section 6 assumes either an explicit instruction or format component in the reward, or at least a separate instruction-following evaluation pack.
3.3 Safety and Privacy
Long, explicit reasoning traces widen the attack and leakage surface. Frontier reasoning models are jailbreakable at high rates; automated attacks approach deterministic success in the studied settings, and success rises with attempt and context budgets (Kassianik and Karbasi, 2025; Zhou et al., 2025a; Jiang et al., 2025). Many-shot evaluations show the same scaling. Targeted defenses tuned specifically to that regime can sharply reduce measured attack success in controlled tests, but these gains are brittle across models and budgets and can impose utility costs (Ackerman and Panickssery, 2025). Richer CoT also increases privacy risk: attributes, confidential prompts, and dataset contents can be reconstructed more easily as intermediate reasoning grows (Green et al., 2025). Safety tuning itself can impose a “safety tax,” degrading math and coding unless staged and balanced carefully (Huang et al., 2025; Zhang et al., 2025b).
Threat Model (Reasoning Traces). Who sees traces: exposure risk is highest for open-weight models or deployments that return full CoT; API-only hidden traces lower user-facing risk, while internal logs can aid auditing. Training vs. inference: RLVR increases trace length and attempt budgets during training and (unless capped) inference, which can amplify leakage/jailbreak surfaces. What “risk increase” means: higher probability that sensitive prompt/data fragments or unsafe behaviors appear in controllable intermediate text. When traces help: hidden/internal traces improve monitorability and post-hoc safety audits, so the goal is controlled exposure, not blanket removal.
Takeaway:
These taxes distort reported gains predictably: overconfident hallucinations inflate apparent utility, instruction drift undermines deployability, and expanded attack surfaces raise real-world risk. The answer is not to abandon RLVR but to co-optimize for reliability: compose rewards so correctness, grounding, and abstention can all improve; manage variance and difficulty; and make calibration and provenance part of evaluation. We follow this recipe in the next sections to show where conclusions change once the tax is accounted for in section 4 and section 6.
4 Perspective II: Pitfalls in Evaluation: Are We Measuring Real Reasoning Gains Accurately?
Claims about RLVR progress are highly sensitive to sampling budgets, metric design, and dataset hygiene. Budget mismatch, fragile judge pipelines, and calibration drift can overstate “reasoning gains” and complicate transfer beyond math and code. We treat contamination separately in section 5 and focus here on budgeting, metric robustness, and calibration.
4.1 Budget Parity and Saturation
A recurring pattern in the literature is to report pass@k for RLVR models while holding base models to much smaller budgets (e.g., pass@1 or pass@5). In such settings, measured improvements often reflect extra search rather than a better policy. Prior work stresses matched budgets and decontaminated baselines for fair comparison (Wen et al., 2025a). Small or underspecified benchmarks inflate estimator variance and reduce statistical power, making single-run point estimates unstable; averaging across multiple randomized trials materially improves estimate reliability (Mu et al., 2025). Under parity controls (e.g., SoberScore, a standardized, matched-budget, multi-seed evaluation that fixes decoding/setup and reports mean $\pm$ std), several prominent gaps shrink or disappear (Hochlehnert et al., 2025). In practice, evaluations should match $k$ across base and RLVR, plot saturation curves (accuracy vs. $k$, optionally summarized by area under the curve), and disclose decoding budgets and parameters.
4.2 Metric Fragility and LLM Judges
While RLVR often targets verifiable domains (math/coding), many evaluation targets either lack programmatic verifiers or have only partial/weak ones (e.g., safety/refusal appropriateness, long-form coherence and rationale quality, multi-turn task success, or fuzzy information extraction). The evaluation often relies on LLM judges. LLM-judge pipelines offer convenience but can be brittle: small changes (seed choice, $k$, dataset versioning, instruction placement, option order, or even tensor-parallel settings) produce swings comparable to reported gains (Sun et al., 2025a). Judges are also manipulable (Zhao et al., 2025). Beyond judges, seemingly modest metric or setup shifts can alter conclusions about “stable reasoning” (Liu et al., 2024a). Where programmatic verifiers exist, they should be preferred. When judges are unavoidable, prompt/format perturbations and adversarial probes should be part of the protocol, with robustness deltas reported, inter-judge agreement documented, and all templates/configurations released to reduce hidden degrees of freedom.
4.3 Calibration Drift and Overconfidence
RL optimization often sharpens the output distribution. Top-1 accuracy can rise even as calibration worsens, increasing brittleness under distribution shift (Hochlehnert et al., 2025). Combined with the volatility noted above (Sun et al., 2025a), some apparent “wins” likely reflect confident exploitation of evaluation quirks rather than robust reasoning. Evaluations should therefore track expected calibration error (ECE), output entropy, and refusal/abstention alongside accuracy, and consider early stopping or annealing when calibration degrades even if reward continues to increase. Unless otherwise stated, numeric confidence is obtained from logit-derived probabilities under the model, computed from the normalized likelihood of the extracted final answer under the same decoding setup.
Table 2: Reported scores vs. standardized evaluation (avg@32 of pass@1, estimated by averaging 32 independent single-sample decodes) with matched decoding budgets and a shared verifier/prompt family. $\Delta$ denotes Reported - Standardized. See Table 4 in the Appendix for comprehensive standardized results of recent RLVR models.
Model (checkpoint) Benchmark Reported (setup) Standardized Eval $\Delta$
AIME-24 48.13 (pass@1; Liu et al. 2025a) 45.62 +2.51+2.51
AIME-25 33.33 (pass@1; Liu et al. 2025a) 33.85 0.52-0.52
Nemotron-Research- AMC-23 79.29 (pass@1; Liu et al. 2025a) 85.70 6.41-6.41
Reasoning-Qwen-1.5B v1 Math 91.89 (pass@1; Liu et al. 2025a) 92.01 0.12-0.12
Minerva 47.98 (pass@1; Liu et al. 2025a) 39.27 +8.71+8.71
Olympiad 60.22 (pass@1; Liu et al. 2025a) 64.56 4.34-4.34
AceReason-Nemotron-14B AIME-24 78.60 (avg@64; Yang et al. 2025) 77.29 +1.31+1.31
AIME-25 67.40 (avg@64; Yang et al. 2025) 66.04 +1.36+1.36
DAPO-Qwen-32B AIME-24 50.00 (accuracy; Yu et al. 2025) 51.56 1.56-1.56
Open-RS3-1.5B AIME-24 46.70 (pass@1; Dang and Ngo 2025) 30.94 +15.76+15.76
STILL-3-1.5B AIME-24 39.33 (avg@64; T=0.6T{=}0.6, top-p=0.95p{=}0.95; Min et al. 2025) 31.46 +7.87+7.87
DeepScaleR-1.5B AIME-24 43.10 (pass@1; Luo et al. 2025) 38.54 +4.56+4.56
Polaris-7B-Preview AIME-24 72.60 (avg@32; NLP and collaborators 2025) 66.46 +6.14+6.14
AMC-23 89.00 (avg@8; NLP and collaborators 2025) 93.59 4.59-4.59
4.4 Reported vs. Reproduced: a Gap Analysis
To make these issues concrete, we compared widely cited checkpoints to budget parity controlled runs using the same verifier, matched decoding budgets, and aligned dataset versions.
Important: some sources report pass@1 while others report avg@64 or alternative aggregation (e.g., majority vote) under different decoding settings; in those cases, $\Delta$ conflates metric and budget differences with genuine policy differences. We therefore treat Table 2 as evidence of evaluation sensitivity. Furthermore, the gaps in Table 2 are exacerbated by three main factors:
Sampling budgets.
Multi-sample reporting (e.g., avg@64 with T=0.6T{=}0.6, top-p=0.95p{=}0.95 for STILL-3-1.5B) inflates scores relative to single-shot pass@1 under the same verifier (Min et al., 2025). In our standardized runs, this manifests as +7.87+7.87 for STILL and +6.14+6.14 for Polaris-7B on AIME-24 (NLP and collaborators, 2025).
Template and version drift.
Changes in prompt templates and dataset slices (e.g., AIME curation, AMC answer formats) shift accuracy by several points, with exact hashes frequently missing from model cards (NLP and collaborators, 2025; Liu and contributors, 2025). This helps explain mixed signs for Nemotron-1.5B v1 (Minerva +8.71+8.71 vs. AMC-23 6.41-6.41).
Metric and decision rules.
Sources sometimes mix “accuracy,” pass@k, avg@k (mean of pass@1 across draws), and maj@k (majority vote), or substitute judges where verifiers exist. Once we normalize to a single verifier/spec, several deltas compress (e.g., DAPO-32B on AIME-24: 1.56-1.56(Sun et al., 2025a; Zhao et al., 2025).
Finally, small-set variance matters: on AIME-24/25, seed and decode settings alone can yield $\pm$3–5 percentage points. Confidence intervals and saturation curves are needed to separate noise from effect (Mu et al., 2025; Hochlehnert et al., 2025).
Why this matters
Under unmatched budgets, unprobed judges, and untracked calibration, distributional sharpening can masquerade as capability expansion. Clean, parity-controlled evaluation often shrinks or flips celebrated gains. Hence, our minimum standard (budget parity and saturation curves, variance disclosure, judge robustness, calibration metrics, and a contamination audit in section 5) is necessary to distinguish real gains from artifacts.
5 Perspective III: Data Contamination
Data provenance is a first-order confound: if pretraining or RL data overlap evaluation sets, measured “reasoning” may reflect memorization. Contamination-aware evaluations in coding (Liu et al., 2024b) and multi-domain leaderboards (Liang et al., 2023) already recommend strict provenance checks; our partial-prompt audit provides direct evidence in math. Partial-prompt reconstruction is a high-precision but not exhaustive signal of overlap: we treat it as one useful probe to be triangulated with lexical and fuzzy matching and with fresh test sets.
Table 3: Partial-prompt contamination summary. The Math block (left) reports answer-match accuracy at an 80% prefix on two legacy sets vs. the fresh AIME-2025. The SimpleQA block (right) is a non-math control, where Qwen’s advantage largely attenuates. Full results, including ACC@60/40 and ROUGE-L/EM for all models and datasets, are in Tables 610 in the Appendix.
Model Math (ACC@80) SimpleQA
MATH-500 AIME-2025 R@80 EM@80
Llama-3-1.8B 2.8 0.0 37.11 19.86
Qwen2.5-Math-7B 58.0 0.0 29.34 12.47
Qwen2.5-32B 60.0 3.3 41.06 23.09
Qwen3-14B-Base 58.2 0.0 37.24 19.40
Probe design (partial-prompt completion).
Following Wu and others (2025), we reveal only the first $x\%$ of each problem ($x\in\{80,60,40\}$), greedily decode the suffix, and compute ROUGE-L@$x$, EM@$x$, and ACC@$x$. High EM/ACC under large prefixes indicates the model can reconstruct the hidden tail and final answer verbatim.
Case study: Qwen vs. Llama.
As summarized in Table 3, Qwen2.5-Math-7B and Qwen3-14B-Base achieve high ACC@80 on legacy math sets (around 58–60% on MATH-500) yet collapse on AIME-2025 (0–3%). Llama-3-1.8B remains near zero across the same sets, consistent with clean evaluation. This pattern (strong on older math sets, absent on the newest release) is consistent with substantial overlap between legacy math benchmarks and the training corpora of some Qwen-family models, and with comparatively cleaner evaluation for Llama. (See Appendix Tables 69 for ACC/ROUGE/EM at 80/60/40% prefixes.) Similar contamination-aware stance and methodology are advocated in code by Liu et al. (2024b) and in broader evaluations by Liang et al. (2023).
SimpleQA control.
On a non-math control (SimpleQA) there is no systematic Qwen advantage: for example, Qwen2.5-32B attains 41.06/23.09 (R@80/EM@80) versus Llama-3-1.8B at 37.11/19.86 (Appendix Table 10). The attenuation on an unseen domain supports the interpretation that elevated partial-prompt math scores on legacy sets reflect contamination rather than a general suffix-reconstruction ability. We use SimpleQA as a non-math control to check that strong tail-reconstruction behavior does not appear uniformly across domains. While suffix reconstruction may be easier for templated math than for open-domain QA, the absence of a consistent advantage on SimpleQA suggests that the large Qwen deltas on legacy math are not explained by a domain-general completion skill alone.
Controls and implications.
We (i) treat contaminated sets as probes of memorization, not reasoning; (ii) prioritize uncontaminated or freshly released test sets; and (iii) publish prompts/seeds/configurations to enable third-party screening. When we re-evaluate RLVR vs. base models under these controls, widely cited gaps shrink or flip. This directly supports our thesis: contamination can make modest distributional sharpening appear as frontier expansion.
6 The Tax-Aware Minimum Standard for RLVR
Our results in section 4 and section 5 show that three pieces of evaluation controls materially change how RLVR gains appear: parity-controlled budgeting, calibration and attempt tracking, and contamination audits. Alongside synthesized prior work on judge robustness and safety evaluations, we package these into a tax-aware minimum standard for reporting RLVR results. The standard is intentionally narrow and measurement-focused. It does not prescribe a particular RL algorithm. Instead, it specifies the controls that must be in place before headline reasoning gains are treated as reliable.
(1) Budget parity and saturation curves.
First, RLVR models and their baselines must be evaluated under matched sampling budgets. In our gap analysis, several celebrated improvements shrank or disappeared once we fixed the verifier, prompt family, decoding parameters, and number of samples per item (Table 2). Reporting pass@k for the RLVR model and pass@1 for the base model, or silently changing temperature, top-$p$, or stopping rules, makes it impossible to tell how much of the gain comes from extra search rather than a better policy.
Under the minimum standard, any claim that RLVR improves a metric on a given benchmark must include: (i) the exact sampling budget for both base and RLVR models, (ii) a saturation curve that plots accuracy as a function of $k$ under a shared decoding setup, and (iii) at least three seeds with mean and confidence intervals or standard deviations (Mu et al., 2025; Hochlehnert et al., 2025). We recommend summarizing performance by area under the saturation curve in addition to a single pass@k point, since this is less sensitive to one particular choice of $k$ and reveals whether RLVR shifts the whole budget-performance frontier or only improves at very large $k$.
(2) Calibration, abstention, and judge robustness.
Second, evaluations must track calibration and abstention, not only accuracy. section 3 showed that RLVR often reduces refusals and increases stated confidence, which yields more attempted items but also more confident errors. Our factual QA control (Table 1) separates shared accuracy from newly attempted tail items and measures expected calibration error (ECE) on shared items. This reveals a distinct failure mode: headline scores rise because the model stops saying “I do not know” and starts answering everything, while accuracy on overlapping items barely moves.
Under the minimum standard, each reported accuracy figure must be accompanied by: (i) refusal or “not attempted” rates, (ii) shared accuracy on the intersection of items both models attempted, and (iii) a calibration metric such as ECE computed from the model’s confidence scores (Hochlehnert et al., 2025; Leng et al., 2024). For settings that rely on LLM-as-a-judge rather than programmatic verifiers, we also require at least one judge stress test: the same outputs scored under several prompt templates or instruction orderings, with the spread in scores reported (Zhao et al., 2025; Sun et al., 2025a; Yang et al., 2026). This does not remove judge fragility, but it makes visible how sensitive a claimed RLVR gain is to small changes in the judge pipeline. Together, these measurements ensure that gains are not driven primarily by attempt inflation, miscalibration, or a brittle judge configuration.
(3) Contamination audits and data hygiene.
Third, RLVR claims must be supported by explicit data provenance checks. section 5 used partial-prompt completion to show that some math benchmarks are heavily memorized by Qwen families: the model can reconstruct the hidden tail and final answer when given 80 percent of the problem on legacy sets, but not on fresh ones such as AIME-2025. Without such checks, apparent reasoning gains can simply reflect better recall of training data (Liu et al., 2024b; Liang et al., 2023; Wu and others, 2025).
Under the minimum standard, any benchmark used to support an RLVR claim must be accompanied by: (i) a contamination screen that combines fuzzy or lexical matching against pretraining and fine-tuning corpora (where available) and partial-prompt probes at several prefix lengths, and (ii) at least one clean held-out set that shows no evidence of tail reconstruction even at large prefixes. For benchmarks that do show contamination, we treat them as probes of memorization rather than as primary evidence of reasoning and clearly label them as such. We also require that dataset versions, prompt templates, and filtering rules be released or described precisely enough for others to re-run the audit.
Summary.
The tax-aware minimum standard asks for three things before treating RLVR gains as robust: matched budgets with saturation curves and variance disclosure, calibration and abstention metrics with at least one judge stress test when judges are used, and a contamination audit with at least one clean held-out benchmark. Taken together, these controls provide a simple, concrete bar that future RLVR work can meet without specifying a particular training recipe, and that can be reused for SFT, RLHF, and test-time compute evaluations.
7 Conclusion
Our central position is that RLVR is effective and deployable for verifiable domains, but headline “reasoning” gains should be treated as provisional unless a small core of tax-aware controls (budget parity, calibration/abstention tracking, robust evaluation with at least one judge stress test, and a simple contamination audit) is enforced. RLVR delivers real gains on verifiable tasks, but the field often over-indexes on headline accuracy while under-weighting taxes and measurement. Under budget parity controls and calibration tracking, several celebrated “reasoning gains” shrink, suggesting that part of the progress reflects distributional sharpening rather than durable expansion.
Broader relevance. The same pressures arise in SFT, RLHF, and test-time compute (e.g., CoT and maj@k). This tax-aware minimum standard applies unchanged beyond RLVR.
Why focus on RLVR. We focus on RLVR because verifiers make these trade-offs measurable and make RLVR a useful proving ground. As a result, it is a useful proving ground for reliability methods that transfer to broader LLM training and deployment. We hope this synthesis helps the community separate genuine capability expansion from artifacts of budgets, metrics, and data, and encourages tax-aware training and reporting across RLVR, SFT, and RLHF.
Limitations
Our goal is measurement clarity rather than a complete theory of when RLVR yields capability expansion, so our evidence has clear boundaries. First, our empirical analyses focus on verifiable domains (math, code, structured QA) using open checkpoints and public benchmarks; findings may not transfer directly to agentic, multimodal, or tool-use settings where programmatic verification is harder. Second, our attempt inflation and calibration analyses depend on operationalizing "attempts" (e.g., extractable answers versus explicit refusals); different extraction rules shift absolute rates, so we emphasize shared-item comparisons to reduce sensitivity. Third, our partial-prompt contamination probe is a high-precision indicator of tail reconstruction on some datasets but is not exhaustive: it can miss semantic overlap and does not quantify training-set membership. We treat it as one triangulation signal alongside provenance checks and fresh held-out sets. We omit approximate nearest-neighbor or fuzzy-hash matching against training corpora and shard-level provenance analysis, which would strengthen contamination claims but require additional infrastructure and training-data access. Finally, we exclude RLVR reward-component ablations (e.g., explicit refusal, grounding, or calibration terms) because they require additional training runs; we view them as promising next steps.
Ethical Considerations
This paper is primarily about evaluation and reporting practices, but it touches on safety and privacy. Longer reasoning traces can increase the chance of exposing sensitive strings or unsafe intermediate content. To mitigate risk, we restrict experiments to public benchmarks and report aggregate statistics rather than releasing jailbreak payloads or automated attack scripts. We also highlight calibration, abstention, and provenance as first-class reporting targets because inflated headline accuracy can create misleading impressions of real-world reliability. Finally, our recommendations can increase evaluation cost (e.g., multi-seed, budget-matched saturation curves); we view this as a necessary trade-off for measurement integrity when results are used to justify deployment.
Acknowledgments
This work was supported in part by Google Cloud credits through the Gemini Academic Program, which enabled large-scale evaluation and experimentation. We also thank DSO National Laboratories (Singapore) for support on knowledge-focused reasoning research. The views and conclusions expressed in this paper are those of the authors and do not necessarily reflect the official policies of the sponsors.
Appendix A Appendix
Conventions and metrics.
pass@k = probability at least one of $k$ samples is correct; avg@k = mean pass@1 across $k$ draws; maj@k = majority vote over $k$; ECE = expected calibration error. Arrows ($\uparrow$$\downarrow$) indicate whether higher or lower is better.
A.1 Supplementary Figure for the Introduction
This appendix subsection contains Figure 2, which shows the RLVR activity trend alongside AIME performance.
Figure 2: Monthly RLVR activity vs. AIME performance (time span: May 2024–June 2025). Left axis: count of pages per month whose title or abstract contains “RLVR” or “reinforcement learning with verifiable rewards” (Google Scholar and arXiv via SerpAPI). Right axis: pass@1 (%) on AIME-24 (circles) and AIME-25 (squares) for selected models; labels show model names.
A.2 Potential Criticisms and Responses
On capability expansion.
RLVR can expand capability under specific regimes (e.g., prolonged RL with resets and KL control; unlikeliness-based shaping; domain-targeted RLVR with sparse pretraining) (Liu et al., 2025a; He et al., 2025a; Cheng et al., 2025). Our claim is not negation but measurement: without tax-aware training and parity-controlled evaluation, expansion is often overstated. The protocol in section 6 is compatible with these positive regimes while curbing overclaiming.
On “it’s just data.”
Data quality and curricula matter (Chen et al., 2025a; Wen et al., 2025a), but reward/metric design independently shapes failure modes (determinacy/overconfidence under judge rewards; terse/off-topic outputs under naive factuality; calibration drift) (Chen et al., 2025b; Leng et al., 2024; Hochlehnert et al., 2025). We treat both data and objectives as first-class knobs (see section 6).
On budgeted pass@k.
We use pass@k for practical relevance, but ask for matched budgets and saturation curves. Under these controls, headline gaps often shrink or flip (Hochlehnert et al., 2025). Reporting AUC alongside pass@k keeps comparisons fair (section 4).
On safety/privacy risk.
Jailbreak success scales with attempts/context; many-shot tuning can reduce measured rates in controlled settings (Zhou et al., 2025a; Jiang et al., 2025; Ackerman and Panickssery, 2025). Longer CoT increases leakage surfaces (Green et al., 2025; Zhou et al., 2025b). Our protocol caps Best-of-NN/CoT length at eval and co-optimizes abstention/privacy with accuracy (section 6).
On reliance on LLM judges.
When verifiers are unavailable or partial, we use LLM judges with robustness probes (prompt/format perturbations, adversarial tests) and publish templates/configs; judges are manipulable, so robustness deltas are reported (Zhao et al., 2025; Sun et al., 2025a). Prefer verifiers when possible (section 4).
Scope.
Our synthesis centers on verifiable math/code/QA with open-weight models. Agentic/multimodal settings introduce additional privacy/safety channels and may need tailored verifiers and audits; we expect core principles (budget parity, calibration, abstention, contamination screens) to transfer with domain-specific adaptations.
A.3 Standardized Evaluations
We report full results under matched budgets, fixed verifiers, and a shared prompt family. These tables extend section 4 and underlie the saturation-curve argument: once $k$, templates, and dataset versions are controlled, several headline gaps narrow. See Table 4 below for the full standardized benchmark results. All results in this section use the same verifier and prompt template family; we specify $k$, number of seeds, and dataset versions below.
Key observations.
[…]
Table 4: Standardized evaluation across math benchmarks (higher is better). Styling: within each size family, the best score for a given benchmark is bold and the second-best is underlined. Context: Parity-controlled scores used in section 4. Setup: pass@1 estimated by averaging over k=32k{=}32 independent single-sample decodes (avg@32), same verifier and prompt family, 3 seeds. Compute: evaluations ran on a $128{\times}$H100 (96 GB) cluster and consumed $\approx$3,500 GPU-hours.
[…]
Qwen2.5-Math-1.5B 8.33 6.35 44.06 66.67 18.42 30.74 29.10
Qwen2.5-Math-1.5B-Instruct 10.10 8.85 55.08 74.83 29.32 40.00 36.37
DeepSeek-R1-Distill-Qwen-1.5B 31.15 24.06 72.81 85.01 32.18 51.55 49.46
STILL-3-1.5B 31.46 25.00 75.08 86.24 32.77 53.84 50.73
DeepScaleR-1.5B 38.54 30.52 80.86 88.79 36.19 58.95 55.64
Qwen2.5-Math-1.5B-Oat-Zero 20.00 10.00 52.50 74.20 26.84 37.78 36.89
Open-RS1 30.94 22.60 73.05 84.90 29.92 52.82 49.04
Open-RS2 28.96 24.37 73.52 85.06 29.74 52.63 49.05
Open-RS3 30.94 24.79 72.50 84.47 29.11 52.25 49.01
Nemotron-Research-Reasoning-Qwen-1.5B v1 45.62 33.85 85.70 92.01 39.27 64.56 60.17
Nemotron-Research-Reasoning-Qwen-1.5B v2 51.77 32.92 88.83 92.24 39.75 64.69 61.70
Qwen2.5-Math-7B 15.62 6.56 52.81 67.72 15.64 32.44 31.80
Qwen2.5-Math-7B-Instruct 12.19 9.17 58.36 83.21 35.56 41.60 40.01
DeepSeek-R1-Distill-Qwen-7B 53.23 38.96 89.30 93.95 43.07 66.67 64.20
Qwen2.5-Math-7B-Oat-Zero 26.67 6.67 67.50 79.20 32.72 41.78 42.42
Skywork-OR1-7B 66.88 51.15 92.73 96.04 44.03 73.61 70.74
LEAD-7B 51.67 37.19 89.06 93.73 43.11 66.34 63.52
AceReason-Nemotron-7B 65.83 47.19 95.08 95.81 45.35 73.92 70.53
AceReason-Nemotron-1.1-7B 71.56 64.58 93.36 96.73 44.05 76.37 74.44
Polaris-7B-Preview 66.46 51.56 93.59 95.68 44.47 73.65 70.90
Qwen2.5-14B 11.04 7.92 47.19 73.19 22.51 37.01 33.14
Qwen2.5-14B-Instruct 13.65 12.40 58.13 80.28 38.63 43.23 41.05
DeepSeek-R1-Distill-Qwen-14B 67.81 48.33 95.39 95.28 46.43 72.06 70.88
LEAD-14B 64.06 52.29 92.81 95.23 47.52 72.25 70.69
AceReason-Nemotron-14B 77.29 66.04 98.67 96.90 47.73 77.34 77.33
Qwen2.5-32B 15.62 9.17 59.30 76.51 26.42 41.45 38.08
Qwen2.5-32B-Instruct 17.19 14.17 67.66 83.17 40.96 47.85 45.17
DeepSeek-R1-Distill-Qwen-32B 69.06 55.52 95.62 95.74 46.50 72.76 72.53
DAPO-Qwen-32B 51.56 36.98 92.73 80.74 33.07 48.97 57.34
Enigmata-Qwen2.5-32B 61.67 46.88 91.25 93.69 46.32 69.14 68.16
DeepSeek-V3-0324 55.83 43.33 92.50 95.12 48.58 66.91 67.05
DeepSeek-R1-0528 90.00 78.33 99.38 97.80 48.35 76.44 81.72
Qwen3-4B 21.88 17.92 66.95 84.27 38.50 52.38 46.98
Qwen3-4B (Thinking) 72.40 62.60 95.78 96.31 46.44 76.40 74.99
Polaris-4B-Preview 27.29 23.02 71.56 85.89 39.20 59.13 51.02
Polaris-4B-Preview (Thinking) 79.90 77.71 99.45 97.38 47.21 80.38 80.34
Qwen3-8B 25.42 20.21 69.61 84.54 39.81 54.15 48.96
Qwen3-8B (Thinking) 77.29 66.46 95.00 96.86 49.06 77.56 77.04
DeepSeek-R1-0528-Qwen3-8B 75.73 67.29 97.19 96.32 45.85 74.19 76.10
A.4 Sequence Length and Compute Footprint
We summarize generated token lengths by benchmark as a proxy for compute footprint and exposure surface. Longer chains increase latency and may worsen calibration and safety/privacy risks (section 3), so we report typical generation lengths under the same decoding budgets used in Appendix A.3. See Table 5 below for the token-length breakdown by model and benchmark.
[…]
Model AIME-24 AIME-25 AMC-23 Math Minerva Olympiad Avg
Qwen2.5-Math-1.5B 1114 1033 828 648 959 904 914
Qwen2.5-Math-1.5B-Instruct 990 891 801 566 640 810 783
DeepSeek-R1-Distill-Qwen-1.5B 16363 16252 9979 5700 8194 11873 11394
STILL-3-1.5B 13350 13000 7716 4314 5921 9345 8941
DeepScaleR-1.5B 9780 8978 5003 3139 5270 5807 6330
Qwen2.5-Math-1.5B-Oat-Zero 1166 1155 848 626 689 892 896
Open-RS1 13578 13698 7495 4170 5935 8995 8978
Open-RS2 14215 13618 7612 4174 5805 9059 9081
Open-RS3 14394 13606 7884 4233 5688 9047 9142
Nemotron-Research-Reasoning-Qwen-1.5B 7786 7713 6294 5070 6569 6678 6685
Qwen2.5-Math-7B 1201 1152 946 720 1136 950 1018
Qwen2.5-Math-7B-Instruct 1465 1418 983 670 748 1051 1056
DeepSeek-R1-Distill-Qwen-7B 13613 14543 6402 4125 5595 8988 8878
Qwen2.5-Math-7B-Oat-Zero 1032 1178 884 673 696 870 889
Skywork-OR1-7B 15366 17845 8361 5541 8566 11818 11250
LEAD-7B 10838 11573 4863 3111 3692 7054 6855
AceReason-Nemotron-1.1-7B 14331 16502 6672 3835 6676 9060 9513
Polaris-7B-Preview 12564 14389 6538 4313 6125 8681 8768
Qwen2.5-14B 1076 1088 847 591 1209 815 938
Qwen2.5-14B-Instruct 1079 994 850 607 632 844 834
DeepSeek-R1-Distill-Qwen-14B 11295 13389 5735 3781 4919 8042 7860
LEAD-14B 8364 9019 4469 3031 4675 5713 5879
AceReason-Nemotron-14B 13871 16334 7239 4609 7677 10030 9960
Qwen2.5-32B 1165 1055 782 556 1186 800 924
DeepSeek-R1-Distill-Qwen-32B 10979 13012 5826 3652 4663 7924 $76$
DAPO-Qwen-32B 6627 6074 3086 2812 3887 5122 4601
Enigmata-Qwen2.5-32B 13204 15417 10195 6358 13770 12098 11840
DeepSeek-V3-0324 3601 3474 2134 1281 705 2386 2264
DeepSeek-R1-0528 16276 18416 10285 6041 7085 13750 11976
MiMo-7B-Base 21716 18842 15734 6390 13809 13299 14965
MiMo-7B-SFT 13012 14250 7030 4177 6152 8421 8841
MiMo-7B-RL-Zero 18105 18575 9950 6564 6546 12152 11982
Qwen3-4B 8909 6079 2369 1289 812 3217 3779
Qwen3-4B (Thinking) 14750 17813 8290 5076 6682 10606 10536
Polaris-4B-Preview 5582 6148 2637 1293 804 4632 3516
Polaris-4B-Preview (Thinking) 28369 33109 15125 9333 13307 21832 20179
Qwen3-8B 8375 6151 2863 1479 662 3624 3859
Qwen3-8B (Thinking) 14764 18296 8756 5391 7105 11384 10949
DeepSeek-R1-0528-Qwen3-8B 20970 22742 12895 7640 9329 16108 14947
A.5 Data Contamination Probes: Math
We use partial-prompt completion to detect tail reconstruction: reveal the first $x\%$ of the problem ($x\in\{80,60,40\}$), greedily decode the remainder, and score ACC/EM/ROUGE-L at that prefix. High scores at large prefixes indicate memorization of legacy sets rather than reasoning. Tables here expand section 5. Accuracy results are reported in Tables 6 and 7; text-overlap and exact-match proxies are reported in Tables 8 and 9.
Table 6: Accuracy of Qwen3 checkpoints on five mathematics benchmarks when each model receives only the first $x\%$ of the question ($x=80,60,40$) and must greedily complete the remainder. Columns “ACC ($x\%$)” report the average accuracy at that prefix length. Takeaway: Qwen3 variants achieve high ACC@80 on legacy MATH-500/AMC-23 but collapse on AIME-2025, consistent with contamination in older sets.
Model Dataset ACC (80%) ACC (60%) ACC (40%) MATH-500 26.40 14.40 5.40 AMC-23 20.00 5.00 0.00 Qwen3-0.6B-Base AIME 2024 3.33 0.00 0.00 AIME 2025 0.00 0.00 0.00 Minerva-Math 1.10 0.00 0.00 MATH-500 33.80 18.40 9.00 AMC-23 20.00 15.00 0.00 Qwen3-1.7B-Base AIME 2024 0.00 0.00 0.00 AIME 2025 3.33 3.33 0.00 Minerva-Math 2.94 2.21 0.74 MATH-500 43.20 28.20 15.00 AMC-23 37.50 37.50 17.50 Qwen3-4B-Base AIME 2024 10.00 6.67 10.00 AIME 2025 3.33 0.00 0.00 Minerva-Math 4.78 2.31 0.74 MATH-500 52.00 35.60 24.80 AMC-23 35.00 25.00 27.50 Qwen3-8B-Base AIME 2024 13.33 6.67 10.00 AIME 2025 3.33 3.33 0.00 Minerva-Math 5.88 3.31 1.84 MATH-500 58.20 44.80 30.20 AMC-23 47.50 37.50 32.50 Qwen3-14B-Base AIME 2024 10.00 16.67 16.67 AIME 2025 0.00 0.00 0.00 Minerva-Math 5.88 2.21 2.94 MATH-500 62.40 51.40 37.80 AMC-23 40.00 25.00 25.00 Qwen3-30B-A3B AIME 2024 13.33 23.33 13.33 AIME 2025 3.33 0.00 0.00 Minerva-Math 4.78 3.21 1.84
Table 7: Accuracy of Qwen2.5 and Llama-3 checkpoints on five mathematics benchmarks when each model receives only the first $x\%$ of the question ($x=80,60,40$) and must greedily complete the remainder. Takeaway: Qwen2.5 shows strong tail reconstruction on legacy math; Llama-3-1.8B remains near zero across sets, reinforcing the contamination interpretation.
Model Dataset ACC (80%) ACC (60%) ACC (40%) MATH-500 58.00 44.40 29.20 AMC-23 52.50 42.50 32.50 Qwen2.5-Math-7B AIME 2024 16.67 20.00 16.67 AIME 2025 0.00 0.00 0.00 Minerva-Math 7.72 4.41 2.94 MATH-500 44.80 27.80 15.00 AMC-23 27.50 27.50 22.50 Qwen2.5-7B AIME 2024 6.67 0.00 3.33 AIME 2025 3.33 3.33 0.00 Minerva-Math 6.62 4.78 1.47 MATH-500 44.80 28.60 14.00 AMC-23 40.00 15.00 12.50 Qwen2.5-7B-Instruct AIME 2024 3.33 0.00 0.00 AIME 2025 3.33 0.00 0.00 Minerva-Math 10.29 5.51 3.31 MATH-500 50.60 35.80 21.00 AMC-23 40.00 27.50 27.50 Qwen2.5-14B AIME 2024 10.00 3.33 6.67 AIME 2025 3.33 0.00 0.00 Minerva-Math 8.46 6.25 2.21 MATH-500 60.00 47.80 32.00 AMC-23 52.50 45.00 42.50 Qwen2.5-32B AIME 2024 16.67 13.33 10.00 AIME 2025 3.33 0.00 6.67 Minerva-Math 9.93 4.41 2.94 MATH-500 2.80 2.40 2.00 AMC-23 0.00 0.00 0.00 Llama-3-1.8B AIME 2024 0.00 0.00 0.00 AIME 2025 0.00 0.00 0.00 Minerva-Math 1.84 0.37 0.00
Table 8: RougeL and exact-match (EM) scores for Qwen3.0 checkpoints on five mathematics benchmarks when each model receives only the first $x\%$ of the question ($x=80,60,40$) and must greedily complete the remainder. Columns “R@x” and “EM@x” report the average ROUGE-L and EM for that prefix length. Takeaway: Elevated ROUGE-L/EM at large prefixes on legacy sets (but not on AIME-2025) indicates suffix reconstruction rather than emergent reasoning.
Model Dataset R@80 EM@80 R@60 EM@60 R@40 EM@40
MATH-500 53.31 20.80 45.90 6.40 39.33 1.00
AMC-23 52.26 7.50 40.39 0.00 33.92 0.00
Qwen3-0.6B-Base AIME 2024 54.85 20.00 27.79 0.00 24.76 0.00
AIME 2025 54.69 10.00 33.47 0.00 29.16 0.00
Minerva-Math 31.41 1.47 29.17 0.00 24.86 0.00
MATH-500 55.86 24.60 49.05 10.20 41.91 2.40
AMC-23 64.35 30.00 52.57 20.00 44.95 12.50
Qwen3-1.7B-Base AIME 2024 56.55 26.67 36.48 6.67 36.56 6.67
AIME 2025 53.90 16.67 39.40 6.00 31.09 0.00
Minerva-Math 33.99 3.31 30.32 0.00 27.28 0.00
MATH-500 65.04 37.40 55.98 18.80 46.44 7.00
AMC-23 73.24 52.50 68.56 45.00 63.89 35.00
Qwen3-4B-Base AIME 2024 68.98 43.33 53.22 30.00 53.22 26.67
AIME 2025 55.34 10.00 39.29 0.00 27.72 0.00
Minerva-Math 34.13 2.94 32.19 0.37 28.18 0.00
MATH-500 70.98 47.20 63.69 29.60 53.84 15.20
AMC-23 78.23 57.50 69.09 52.50 69.73 47.50
Qwen3-8B-Base AIME 2024 77.28 60.00 55.76 33.33 57.85 30.00
AIME 2025 51.97 10.00 38.60 0.00 32.21 0.00
Minerva-Math 35.87 2.57 33.41 0.00 28.52 0.00
MATH-500 74.93 55.80 70.08 39.80 60.23 23.80
AMC-23 75.32 55.00 73.20 60.00 75.52 52.50
Qwen3-14B-Base AIME 2024 72.31 50.00 61.48 40.00 60.54 40.00
AIME 2025 56.07 10.00 36.41 0.00 32.46 0.00
Minerva-Math 38.12 4.41 34.11 0.74 29.72 0.00
MATH-500 80.12 62.40 75.03 45.60 64.93 33.00
AMC-23 80.40 60.00 73.24 55.00 77.21 55.00
Qwen3-30B-A3B AIME 2024 74.26 53.33 66.53 40.00 62.90 33.33
AIME 2025 51.57 6.67 43.18 0.00 30.55 0.00
Minerva-Math 37.01 3.68 34.96 1.10 29.22 0.00
Table 9: RougeL and exact-match (EM) scores for Qwen2.5 and Llama-3 on five mathematics benchmarks when each model receives only the first $x\%$ of the question ($x=80,60,40$) and must greedily complete the remainder. Takeaway: Qwen2.5 shows strong reconstruction on legacy math; Llama-3 remains low.
Model Dataset R@80 EM@80 R@60 EM@60 R@40 EM@40
MATH-500 79.63 61.60 70.30 41.20 60.81 25.40
AMC-23 77.35 57.50 71.47 50.00 66.29 42.50
Qwen2.5-Math-7B AIME 2024 69.35 53.33 55.90 30.00 56.53 30.00
AIME 2025 48.20 10.00 36.31 0.00 27.23 0.00
Minerva-Math 33.42 2.94 30.41 0.37 26.88 0.00
MATH-500 66.09 38.20 58.15 18.00 48.34 6.00
AMC-23 69.47 47.50 67.70 45.00 61.08 35.00
Qwen2.5-7B AIME 2024 61.28 30.00 46.19 16.67 46.14 16.67
AIME 2025 54.62 10.00 42.41 0.00 30.56 0.00
Minerva-Math 33.88 2.94 31.13 0.00 27.06 0.00
MATH-500 60.05 28.60 51.10 10.40 43.41 2.60
AMC-23 61.38 30.00 51.22 17.50 42.90 10.00
Qwen2.5-7B-Instr AIME 2024 55.41 20.00 36.28 0.00 30.64 0.00
AIME 2025 58.12 10.00 37.96 0.00 29.81 0.00
Minerva-Math 34.37 2.94 28.50 0.00 25.99 0.00
MATH-500 69.50 45.00 61.20 27.00 51.24 12.80
AMC-23 74.58 47.50 68.20 42.50 67.82 37.50
Qwen2.5-14B AIME 2024 65.42 40.00 49.76 23.33 48.97 23.33
AIME 2025 52.56 6.67 38.02 0.00 32.92 0.00
Minerva-Math 35.64 3.68 31.30 0.00 27.93 0.00
MATH-500 76.98 56.60 71.14 41.20 58.32 25.20
AMC-23 79.18 62.50 74.59 60.00 75.78 55.00
Qwen2.5-32B AIME 2024 73.25 56.67 60.92 40.00 60.47 33.33
AIME 2025 53.67 6.67 40.91 0.00 31.90 0.00
Minerva-Math 34.88 2.94 32.92 0.00 28.32 0.00
MATH-500 46.49 14.00 38.75 3.00 32.19 0.60
AMC-23 40.46 0.00 29.87 0.00 25.28 0.00
Llama-3-1.8B AIME 2024 49.70 16.67 31.45 0.00 27.13 0.00
AIME 2025 50.16 6.67 32.97 0.00 27.18 0.00
Minerva-Math 35.32 1.47 29.17 0.00 26.95 0.00
A.6 Data Contamination Probes: SimpleQA (Control)
To test whether partial-prompt reconstruction reflects a general suffix-completion ability, we repeat the probe on SimpleQA. If math-set effects were general, we would expect similarly high ROUGE-L/EM at large prefixes here. They are not: scores cluster modestly across families. See Table 10 below for the ROUGE-L and exact-match results across prefix lengths.
Table 10: Rouge-L and exact-match (EM) scores on SimpleQA when each model receives only the first $x\%$ of the question ($x=80,60,40$) and must greedily complete the remainder. Takeaway: No consistent Qwen advantage; effects seen on legacy math do not generalize, supporting the contamination interpretation.
Model Dataset R@80 EM@80 R@60 EM@60 R@40 EM@40
Qwen2.5-Math-7B SimpleQA 29.34 12.47 16.19 0.69 13.28 0.00
Qwen2.5-7B SimpleQA 37.61 19.63 19.88 1.85 14.64 0.23
Qwen2.5-7B-Instruct SimpleQA 36.06 17.78 19.76 1.62 15.87 0.46
Qwen2.5-14B SimpleQA 39.97 21.94 21.69 3.23 15.98 0.23
Qwen2.5-32B SimpleQA 41.06 23.09 21.87 3.23 14.78 0.00
Llama-3-1.8B SimpleQA 37.11 19.86 20.24 2.08 14.22 0.00
Qwen3-0.6B-Base SimpleQA 26.45 10.39 16.35 0.46 14.42 0.00
Qwen3-1.7B-Base SimpleQA 27.55 12.47 16.50 0.92 14.97 0.23
Qwen3-4B-Base SimpleQA 33.52 16.40 17.87 1.85 14.93 0.00
Qwen3-8B-Base SimpleQA 33.99 16.40 19.23 1.39 15.39 0.00
Qwen3-14B-Base SimpleQA 37.24 19.40 21.44 2.31 16.65 0.23
Qwen3-30B-A3B-Base SimpleQA 38.80 19.63 22.07 3.00 16.55 0.00
Qwen3-23B-A22B-Instr-2507 SimpleQA 53.88 38.34 34.58 10.39 22.74 1.39